### v1

# ADHint: Adaptive Hints with Difficulty Priors for Reinforcement Learning

Abstract
To combine the advantages of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL), recent methods have integrated “hints” into post-training, which are prefix segments of complete reasoning trajectories, aiming for powerful knowledge expansion and reasoning generalization. However, existing hint-based RL methods typically ignore difficulty when scheduling hint ratios and estimating relative advantages, leading to unstable learning and excessive imitation of off-policy hints. In this work, we propose ADHint, which treats difficulty as a key factor in both hint-ratio schedule and relative-advantage estimation to achieve a better trade-off between exploration and imitation. Specifically, we propose Adaptive Hint with Sample Difficulty Prior, which evaluates each sample’s difficulty under the policy model and accordingly schedules an appropriate hint ratio to guide its rollouts. We also introduce Consistency-based Gradient Modulation and Selective Masking for Hint Preservation to modulate token-level gradients within hints, preventing biased and destructive updates. Additionally, we propose Advantage Estimation with Rollout Difficulty Posterior, which leverages the relative difficulty of rollouts with and without hints to estimate their respective advantages, thereby achieving more balanced updates. Extensive experiments across diverse modalities, model scales, and domains demonstrate that ADHint delivers superior reasoning ability and out-of-distribution generalization, consistently surpassing existing methods in both pass@1 and avg@8. Our code and dataset will be made publicly available upon paper acceptance. *: Equal contribution, 🖂: Corresponding author.
1 Introduction
(a) ADHint w/ annealing hint ratio
(b) ADHint w/ standard advantages
Figure 1: Baseline behavior with existing hint-based designs. “ADHint w/ annealing hint ratio” replaces our sample difficulty prior-based hint-ratio schedule (AH-SDP) with a commonly used time-varying schedule. “ADHint w/ standard advantages” replaces our rollout difficulty posterior-based advantage estimation (AE-RDP) with the standard estimation that pools hint-rollouts and naive-rollouts into a single GRPO group. In each subfigure, the Baseline curve matches the variant in the title, and “Baseline: Hint/Naive” reports statistics from hint- or naive-rollouts only.
Recent advancements in Large Language Models (LLMs) [yang2024qwen2, yang2025qwen3, li2025system, zeng2025glm] and Multimodal Large Language Models (MLLMs) [bai2025qwen3vltechnicalreport, liu2025visual, huang2025vision, wang2025internvl3] highlight the potential to enhance reasoning capabilities by large-scale Reinforcement Learning with Verifiable Rewards (RLVR) [shao2024deepseekmath, guo2025deepseek, yu2025dapo], such as Proximal Policy Optimization (PPO) [schulman2017proximal] and Group Relative Policy Optimization (GRPO) [shao2024deepseekmath, guo2025deepseek]. However, RLVR primarily reinforces existing behaviors rather than instilling novel reasoning abilities [chusft, li2025system]. On the contrary, Supervised Fine-Tuning (SFT) offers a complementary approach to facilitating the incorporation of new knowledge and reasoning patterns [chen2025sft, yan2025learning]. Nevertheless, the SFT models often exhibit limited generalization to out-of-distribution scenarios. These contrasting characteristics of RL and SFT motivate a promising research direction: How can RL and SFT effectively collaborate to enhance both reasoning abilities and out-of-distribution generalization?
To achieve this goal, early methods [yan2025learning, ma2025learning, dong2025rl, fu2025srft] attempt to incorporate on-policy RL with full reasoning trajectories to internalize external knowledge. However, these methods often struggle with insufficient exploration [wu2025templaterlstructuredtemplateguidedreinforcement, xi2025bapostabilizingoffpolicyreinforcement] and even training collapse [liang2025squeezesoakedspongeefficient, cui2025entropymechanismreinforcementlearning]. Recent methods [Huang_2025_ICCV, liu2025uft, zhang2025stephint] introduce hints, which refer to the prefix segments of complete reasoning trajectories, allowing the policy model to explore the continuation rather than reinforcing the model to mimic the entire sequence. Specifically, these methods assign fixed [zhang2025stephint] or time-varying [liu2025uft, huang2025blending] hint ratios to all samples (or only to the hardest samples [Huang_2025_ICCV, liu2025ghpo]), and then ultize hint-rollouts derived from hint-guided reasoning together with naive-rollouts generated without hints for group relative advantage estimation [zhang2025stephint, huang2025blending].
Despite their empirical successes, existing hint-based RL methods are sensitive to sample-difficulty in the hint-ratio schedule and rollout-difficulty in the relative-advantage estimation. Specifically, most previous methods apply a uniform hint ratio to both easy and hard samples for model reasoning, and then generate hint-rollouts with various difficulties. Policy learning from such heterogeneous rollouts could introduce high-variance model optimization, and thus undermine training stability [bengio2009curriculum, wang2025dump, parashar2025curriculum]. As shown in Figure 1(a), the model with annealing hint ratio (denoted as Baseline) meets with training collapse since its entropy rises sharply at the end of training, indicating insufficient absorption of external knowledge and a drift in the model’s own policy. In contrast, when introducing the hint-ratio schedule with sample-difficulty (our method, discussed later), the entropy converges stably so that the model could learn reasoning ability, not superficial patterns in off-policy hints.
As for relative-advantage estimation, most existing approaches pool hint-rollouts and naive-rollouts into a single group so that the relative-advantage is dominated by hint-rollouts, with the model learning to directly imitate the off-policy hint distribution rather than exploring with its own policy under hint guidance. Specifically, externally guided hint-rollouts are typically simpler than naive-rollouts, leading to more trajectories with positive advantages that dominate the update signal and bias the policy toward fitting the off-policy distribution. Additionally, hints from external reasoning models are usually much longer than the policy’s own outputs, further exacerbating the imbalance. As shown in Figure 1(b), the average reward of hint-rollouts increases steadily, whereas that of naive-rollouts collapses, causing the model (denoted as Baseline) to eventually lose the ability to perform inference without hints.
Figure 2: Overall comparison between ADHint and baselines. Each axis corresponds to a dataset–metric pair, where p@1 denotes pass@1, and a@8 denotes avg@8. ADHint achieves the best or highly competitive performance on most metrics.
To address these issues, we propose ADHint (ADaptive Hints with Difficulty Priors for Reinforcement Learning), which incorporates difficulty into both the hint-ratio schedule and the relative-advantage estimation to achieve a principled trade-off between exploration and imitation. For the hint-ratio schedule, we propose Adaptive Hint with Sample Difficulty Prior (AH-SDP), which adaptively determines the hint ratio for each sample based on the difficulty prior estimated from the average reward of its naive-rollouts, keeping the resulting hint-rollouts within a moderate difficulty range to provide stable update signals. Furthermore, to prevent large shifts toward the off-policy distribution that trigger entropy collapse, we propose Consistency-based Gradient Modulation (CGM), which downweights the gradients of hint tokens whose entropy deviates significantly from that of the policy-generated continuation. We also introduce Selective Masking for Hint Preservation (Selective Masking), which filters out erroneous update signals from hint-rollouts with negative advantages. For the relative-advantage estimation, we propose Advantage Estimation with Rollout Difficulty Posterior (AE-RDP), which constructs the rollout difficulty posterior from the average rewards of both naive-rollouts and hint-rollouts, and uses it to estimate advantages. Positive naive-rollouts with greater difficulty receive larger advantages, as they provide valuable learning signals that are better aligned with current policy, whereas negative hint-rollouts with lower difficulty are penalized more heavily.
We conduct extensive experiments across both multimodal and text-only settings, multiple model scales, and diverse benchmarks covering mathematical, multi-disciplinary, and logical reasoning. As shown in Figure 2, ADHint consistently yields substantial gains in pass@1 and avg@8, demonstrating synergistic improvements in both reasoning ability and out-of-distribution generalization. Experiments on the recent, stronger backbone Qwen3-VL-8B-Instruct [yang2025qwen3] further validate the effectiveness of ADHint. Our main contributions are summarized as follows:
  • We reveal that difficulty is a crucial signal for both hint-ratio schedule and relative-advantage estimation, and that neglecting it results in unstable learning and overfitting to the off-policy distribution.
  • We propose ADHint, which explicitly exploits sample difficulty priors and rollout difficulty posteriors for hint-ratio schedule and relative-advantage estimation, leading to a better balance between exploration and imitation.
  • Extensive experiments across diverse settings consistently demonstrate ADHint’s strong performance, delivering significant improvements in pass@1 and avg@8 on multi-domain benchmarks.
2 Related Work
2.1 Multimodal Reinforcement Learning
GRPO [guo2025deepseek, shao2024deepseekmath] removes the critic network from PPO [schulman2017proximal] and estimates advantages via intra-group sampling, which has become a cornerstone for recent reasoning models [yang2025qwen3, zeng2025glm, team2025kimi, team2025longcat, ye2025limo, wang2025reinforcement, tan2025towards]. In parallel, the multimodal community has explored various strategies to enhance MLLM reasoning, including refining reasoning styles [wang2025vl, zheng2025deepeyes, su2025pixel], improving rollout procedures for training robustness [liu2025noisyrollout, wang2025perception], redesigning reward functions [liu2025visionreasoner, liu2025visual], and data augmentation [liang2025modomodomultidomaindatamixtures, li2025truth]. Despite these advances, RL algorithms primarily focus on unlocking the latent potential of pretrained models and often fail to expand their fundamental capability boundary.
2.2 Reinforcement Learning with Off-Policy Data
Recent work explores incorporating off-policy data into RL-based post-training. Methods that directly integrate off-policy data into SFT or RL objectives [yan2025learning, ma2025learning, dong2025rl, fu2025srft] often struggle with the substantial distribution mismatch between off-policy data and the on-policy model. Hint-based methods alleviate this by guiding the model’s reasoning with partial hints extracted from off-policy data while preserving its own exploration. UFT [liu2025uft] and Prefix-RFT [huang2025blending] employ temporally annealed hint ratios, while StepHint [zhang2025stephint] mixes hint-rollouts from multiple hint ratios with naive-rollouts and ground-truth into a single group. Other methods reserve hints only for the hardest samples and determine the hint ratio via uniform sampling [chen2025data], binary search [zhang2025bread], or progressive search [Huang_2025_ICCV, liu2025ghpo]. In this work, we propose ADHint, which explicitly incorporates difficulty into both the hint-ratio schedule and the relative-advantage estimation to address the unstable learning and excessive imitation observed in existing hint-based methods. More detailed discussions are provided in Appendix B.
3 Preliminary
Figure 3: Overview of ADHint. For each sample, ADHint successively infers twice to obtain naive-rollouts and hint-rollouts, where the hint ratio of the latter is scheduled based on the difficulty evaluated from the former. It then leverages the relative difficulty of naive-rollouts and hint-rollouts to estimate their respective advantages, and modulates the gradient of each hint token based on the consistency of entropy.
3.1 Group Relative Policy Optimization
For each query $q$ in the question set $Q$, GRPO samples a group of rollouts $\{o_{1},\dots,o_{G}\}$ from the old policy model $\pi_{\theta_{\text{old}}}$, and computes their respective rewards $\{r_{1},\dots,r_{G}\}$. The current policy model $\pi_{\theta}$ is optimized by maximizing the following objective:
$\mathcal{J}_{\text{GRPO}}(\pi_{\theta})={}\mathbb{E}_{q\sim Q,{\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}}$
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Biggl\{\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i}\;-\beta\,D_{\mathrm{KL}}\bigl(\pi_{\theta}\,\|\,\pi_{\text{ref}}\bigr)\Biggr\}.$
where the standard min and clip operations for avoiding extreme gradients are omitted for clarity, and $D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})$ is a KL constraint term. $\hat{A}_{i}$ is the relative advantage, calculated by normalizing the reward within the group:
$\hat{A}_{i}=\frac{r_{i}-\text{mean}(\{r_{1},\dots,r_{G}\})}{\text{std}(\{r_{1},\dots,r_{G}\})}.$
The gradient for a given $q$ and $o_{i}$ is calculated as:
$\nabla_{\theta}\mathcal{J}^{*}_{\text{GRPO}}(\pi_{\theta})\Big|_{q,o_{i}}=\frac{1}{G\cdot|o_{i}|}$
$\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\hat{A}_{i}\cdot\nabla_{\theta}\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t})).$
3.2 Hint-based On-Policy Learning
Recent works use the initial part of off-policy reasoning trajectories from a teacher model as hints to guide rollouts. Given a query $q$, a hint ratio $w$ specifies the prefix length $h$ copied from the off-policy trajectory. Conditioned on $q$, the policy model $\pi_{\theta}$ then produces two types of rollouts:
  • Naive-rollouts: Generate $n$ rollouts $\{o_{1},\dots,o_{n}\}$ without guidance, each sampled as $o_{i}\sim\pi_{\theta}(\cdot\mid q)$.
  • Hint-rollouts: Generate $m$ rollouts $\{o_{n+1},\dots,o_{n+m}\}$ with hint guidance, each formed as $o_{i}=[\,o_{i,<h}\,;\,o_{i,>h}\,]$, where $o_{i,<h}$ is the hint prefix of length $h$ and $o_{i,>h}$ is the continuation sampled as $o_{i,>h}\sim\pi_{\theta}(\cdot\mid q,o_{i,<h})$.
In standard practice, the n+mn{+}m rollouts are pooled into a single group to compute relative advantages. Notably, since hint tokens ($t\leq h$) are not generated by the policy model, following prior work [yan2025learning], we set $\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})=1$ for all hint tokens in hint-rollouts.
4 Method
(a) ADHint w/o CGM
(b) ADHint w/o Selective Masking
Figure 4: Baseline behavior after removing CGM or Selective Masking. The left subfigure shows the baseline’s entropy (left axis) and response length (right axis). The right subfigure compares the entropy of the baseline and GRPO, where “Baseline: Hint/Naive” is computed on hint- or naive-rollouts only.
4.1 Overview
As shown in Figure 3, ADHint comprises four crucial modules: (1) Adaptive Hint with Sample Difficulty Prior (AH-SDP), where the policy model successively performs two inferences to obtain naive-rollouts and hint-rollouts, and accordingly schedules the hint ratio for the latter based on the sample difficulty prior evaluated from the former; (2) Advantage Estimation with Rollout Difficulty Posterior (AE-RDP), which uses the rollout difficulty posterior to estimate relative advantages for naive-rollouts and hint-rollouts, covering both correct and wrong outcomes; (3) Consistency-based Gradient Modulation (CGM), where the gradient of each hint token is modulated according to the consistency of its entropy distribution with that of the continuation; and (4) Selective Masking for Hint Preservation (Selective Masking), which masks the update signals of hint tokens in hint-rollouts that remain incorrect even with guidance. Additionally, Algorithm 1 summarizes the overall procedure of ADHint.
4.2 Adaptive Hint with Sample Difficulty Prior
Assigning the same hint ratio to every sample in the current update batch [zhang2025stephint, huang2025blending, liu2025uft] introduces a mismatch in the difficulty of hint-rollouts across samples, yielding unstable training signals and reducing learning efficiency. As shown in Figure 1(a), while the entropy drops rapidly in the early stage, it rises sharply in the later phase as the hint ratio becomes low. These observations suggest that the model tends to memorize superficial hint patterns rather than internalize the underlying knowledge.
Unlike existing methods, we use the difficulty prior evaluated from each sample’s naive-rollouts to schedule the hint ratio, thereby keeping hint-rollouts within a moderate-difficulty regime (Figure 5(a)) and providing low-variance update signals at every step. For each query $q$, we define the difficulty score of naive-rollouts as
$\text{Diff}_{\text{N}}\;=\;1-\mathcal{M}(1:n)\;=\;1-\operatorname{mean}(r_{1},\dots,r_{n}).$ (1)
We then compute the hint ratio $w$ using a linear function:
$w\;=\;f(\text{Diff}_{\text{N}})\;=\;(w_{\text{max}}-w_{\text{min}})\cdot\text{Diff}_{\text{N}}+w_{\text{min}}+\sigma.$ (2)
where $w_{\text{max}}$ and $w_{\text{min}}$ denote the maximum and minimum ratio hyperparameters, and $\sigma\sim\mathcal{U}(-R,R)$ is uniformly distributed noise.
Finally, the adaptively computed ratio $w$ determines the hint length $h$, which we then use to generate $m$ hint-rollouts $\{o_{n+1},\dots,o_{n+m}\}$.
Consistency-based Gradient Modulation.
Since off-policy hints can differ substantially from the on-policy model in language style, knowledge structure, and response length, indiscriminately learning from hints risks overfitting to off-policy patterns. As shown in Figure 4(a), the entropies of both naive-rollouts and hint-rollouts collapse sharply, and the response length of naive-rollouts quickly approaches that of hint-rollouts, indicating a policy shift toward the off-policy hints. We mitigate this by measuring the consistency between each hint token and the generated continuation that represents the model’s intrinsic distribution, and using this consistency to modulate the gradients of the hint tokens, thereby avoiding destructive distribution shifts in the policy model.
Specifically, for each hint token in a hint-rollout $o_{i}$, we define the token-level entropy as
$H_{i,t}\;=\;\mathcal{H}\!\big(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\big),$ (3)
and the average entropy of the policy-generated continuation as
$\bar{H}_{i}\;=\;\frac{1}{\,|o_{i,>h}|\,}\sum_{t=h+1}^{|o_{i}|}\mathcal{H}\!\big(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\big).$ (4)
We then scale the gradient of each token by
$k_{i,t}\;=\;\begin{cases}g\!\left(\dfrac{H_{i,t}}{\bar{H}_{i}}\right),&1\leq t\leq h,\\[6.0pt] 1,&t>h,\end{cases}$
where $g(x;\alpha)$ is a symmetric cosine-based schedule centered at $g(1)=1$, with a ramp-up on $\alpha,1]$ and a ramp-down on $1,1/\alpha]$ for $0<\alpha\leq 1$:
$g(x;\alpha)\;=\;\begin{cases}\displaystyle\!\sin\!\Big(\tfrac{\pi}{2}\cdot\tfrac{x-\alpha}{1-\alpha}\Big),&\alpha\leq x\leq 1,\\[8.0pt] \displaystyle\!\cos\!\Big(\tfrac{\pi}{2}\cdot\tfrac{x-1}{(1/\alpha)-1}\Big),&1<x\leq 1/\alpha,\\[8.0pt] 0,&\text{otherwise}.\end{cases}$
Selective Masking for Hint Preservation.
Since the adaptively controlled hint-rollouts remain in a moderate-difficulty regime, some trajectories can yield negative advantages. Applying such negative updates to the hint prefix $o_{i,\leq h}$ is counterproductive because this prefix is assumed to be correct. Penalizing it introduces conflicting gradients and destabilizes learning, often triggering an entropy spike as the model’s predictive distribution becomes overly uncertain (Figure 4(b)). Thus, we selectively mask the gradients of hint tokens in this case:
$k_{i,t}\;=\;\begin{cases}0,&1\leq t\leq h\ \text{and}\ \hat{A}_{i}\leq 0,\\[4.0pt] g\!\left(\dfrac{H_{i,t}}{\bar{H}_{i}}\right),&1\leq t\leq h\ \text{and}\ \hat{A}_{i}>0,\\[6.0pt] 1,&t>h.\end{cases}$ (5)
To unify the notation, we fix $k_{i,t}=1$ for naive-rollouts.
4.3 Advantage Estimation with Rollout Difficulty Posterior
When hint-rollouts and naive-rollouts are pooled into a single group to estimate relative advantages [zhang2025stephint, huang2025blending], the update is often biased toward hint-rollouts, since they typically contain more positive trajectories and longer responses. As shown in Figure 1(b), the average reward of hint-rollouts keeps increasing while the reward of naive-rollouts collapses, indicating overfitting to the hint distribution. In this case, the policy improves “text-completion” behavior but sacrifices genuine reasoning ability.
To address this issue, we introduce the rollout difficulty posterior for advantage estimation, which characterizes the relative difficulty between naive-rollouts and hint-rollouts after both rollouts are completed. Our design follows two intuitions. Since naive-rollouts are typically harder and fully generated by the current policy, their positive trajectories provide more informative signals for policy updates. In contrast, hint-rollouts are relatively easier and thus warrant stronger penalties when incorrect.
Specifically, for each query $q$, we define the difficulty score of naive-rollouts in Eq. 1. And similarly define the difficulty score of hint-rollouts as
$\text{Diff}_{\text{H}}\;=\;1-\mathcal{M}(n+1:)\;=\;1-\operatorname{mean}(r_{n+1},\dots,r_{n+m}).$ (6)
Then the difficulty score of rollout $o_{i}$ is denoted as
$\text{Diff}_{i}\;=\;\begin{cases}\text{Diff}_{\text{N}},&1\leq i\leq n,\\ \text{Diff}_{\text{H}},&n{+}1\leq i\leq n{+}m.\end{cases}$
Let $\hat{A}_{i}$ denote the relative advantage of rollout $o_{i}$ estimated by pooling all n+mn{+}m trajectories. We define $s_{i}$ as a sign indicator:
$s_{i}\;=\;\begin{cases}1,&\hat{A}_{i}>0,\\ -1,&\hat{A}_{i}\leq 0.\end{cases}$
Finally, the relative advantages in ADHint are estimated as
$\tilde{A}_{i}\;=\;\left(\frac{\mathcal{M}(1:n+m)}{1-\text{Diff}_{i}}\right)^{s_{i}}\hat{A}_{i}.$ (7)
4.4 Overall Gradient Computation
Building on the above mechanisms, we compute the gradient for a given query $q$ and rollout $o_{i}$ as:
$\nabla_{\theta}\mathcal{J}^{*}_{\text{ADHint}}(\pi_{\theta})\Big|_{q,o_{i}}=\frac{1}{G\cdot|o_{i}|}$ (8)
$\sum_{t=1}^{|o_{i}|}k_{i,t}\cdot\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\tilde{A}_{i}\cdot\nabla_{\theta}\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t})).$
Notably, to avoid inaccurate difficulty prior evaluation caused by formatting errors, we warm up the policy with GRPO for the first five steps so that the model follows the required response format (Figure 5(f)).
(a) Training reward
(b) Policy entropy
(c) Gradient norm
(d) Response length
(e) Clipping ratio of long responses
(f) Format reward
Figure 5: Training dynamics of key metrics for ADHint and GRPO. “ADHint: Hint/Naive” reports metrics computed on hint- or naive-rollouts only, while “ADHint: Overall” aggregates both rollout types. ADHint maintains stable learning and yields concurrent improvements in reasoning generalization and knowledge expansion.
5 Experiments
Method Math Reasoning Multi-Discipline Logical Avg.
MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista
[…]
SFT 51.9/61.3 31.2/54.2 32.5/58.5 28.6/46.3 17.1/29.4 18.1/37.8 29.9/47.9
SFT $\rightarrow$ RL 53.6/63.3 37.4/54.1 38.3/60.6 30.6/47.7 18.7/30.9 18.1/39.9 32.9/49.4
GRPO [shao2024deepseekmath] 63.9/63.1 43.1/45.0 59.0/59.0 47.5/49.1 30.7/30.1 43.8/38.9 48.0/47.5
LUFFY [yan2025learning] 41.2/45.6 31.7/35.0 42.3/46.1 34.3/44.6 20.3/23.1 35.4/38.3 34.2/38.8
StepHint [zhang2025stephint] 34.9/44.6 28.7/32.5 32.0/40.2 18.3/35.7 19.8/26.0 12.5/27.8 24.4/34.5
HintGRPO [Huang_2025_ICCV] 56.7/52.5 37.4/39.7 42.0/40.5 44.4/43.0 27.5/25.1 35.7/33.7 40.6/39.1
GHPO [liu2025ghpo] 57.3/55.2 35.6/35.4 41.0/39.1 44.4/42.6 27.0/26.3 35.3/33.9 40.1/38.8
[…]
SFT 59.0/68.7 40.1/64.0 44.1/70.1 31.2/51.0 20.6/35.6 23.7/46.0 36.5/55.9
SFT $\rightarrow$ RL 64.7/71.1 56.4/67.2 54.1/72.5 35.9/54.2 25.8/38.3 28.8/47.8 44.3/58.5
GRPO [shao2024deepseekmath] 73.4/73.4 55.3/59.2 69.8/68.8 55.3/55.6 40.4/40.0 44.9/46.4 56.5/57.2
LUFFY [yan2025learning] 63.6/61.2 34.6/38.1 52.2/54.8 44.6/47.4 30.6/29.7 40.2/39.2 44.3/45.1
StepHint [zhang2025stephint] 47.2/49.5 27.7/31.4 39.8/41.1 18.6/27.9 9.5/19.3 10.5/22.7 25.6/32.0
HintGRPO [Huang_2025_ICCV] 65.9/64.3 45.9/47.5 42.9/44.7 40.6/39.3 30.2/28.7 36.4/34.7 43.7/43.2
GHPO [liu2025ghpo] 66.1/63.9 52.3/48.9 63.0/52.8 51.7/50.0 35.3/32.3 42.4/40.5 51.8/48.1
[…]
Table 1: Performance comparison of ADHint and baselines on Qwen2.5-VL across six benchmarks spanning three domains. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value). The upper block uses Qwen2.5-VL-3B as the base MLLM, while the lower block uses Qwen2.5-VL-7B. The best and runner-up results are highlighted in bold and underlined, respectively.
5.1 Settings
Training data.
We conduct extensive experiments across both MLLMs and LLMs to validate ADHint’s effectiveness. For MLLMs, we construct our training data from the ViRL39K dataset [wang2025vl], which spans multiple domains, including STEM, chart reasoning, document reasoning, and spatial reasoning. We use Qwen3-VL-30B-A3B-Thinking and Qwen3-VL-235B-A22B-Thinking [bai2025qwen3vltechnicalreport] to sample hints, and remove examples that are considered excessively simple for Qwen2.5-VL-7B-Instruct [bai2025qwen2], resulting in a final set of 17K training queries, termed ViRL-Hint17k. For LLMs, we use 21K examples sampled from public training data [yan2025learning], with hints generated by DeepSeek-R1 [guo2025deepseek]. More details are provided in Appendix C.1.
Benchmarks.
For MLLM evaluation, we use six benchmarks spanning diverse domains, including MathVista [lu2023mathvista], MathVerse [zhang2024mathverse], and We-Math [qiao2024we] for math reasoning; MMMU [yue2024mmmu] and MMMU-Pro [yue2024mmmupro] for multi-discipline reasoning; and LogicVista [xiao2024logicvista] for logical reasoning. For LLM evaluation, we use AIME24, AIME25, AMC [li2024numinamath], Minerva [lewkowycz2022solving], OlympiadBench [he2024olympiadbench], and MATH500 [hendrycks2measuring]. Further details are provided in Appendix C.2.
Evaluation Metrics.
For MLLM evaluation, we report pass@1 and avg@8 for each benchmark. Pass@1 measures the accuracy of a single rollout per sample under greedy decoding, reflecting the model’s out-of-distribution reasoning generalization. Avg@8 computes the average accuracy over eight rollouts per sample at temperature 1.0, reflecting the model’s robustness and mastery of the underlying knowledge. For LLM evaluation, we report either pass@1 or avg@32 depending on the benchmark to mitigate potential variance introduced by the small sizes of some datasets.
Implementation Details.
For MLLM training, we use Qwen2.5-VL-7B-Instruct, Qwen2.5-VL-3B-Instruct, and Qwen3-VL-8B-Instruct, and train them for 250 steps (about 2 epochs). We set the learning rate to $1\times 10^{-6}$, the rollout batch size to 128, and the maximum generation length to 8K. For LLM training, we use Qwen2.5-Math-7B [yang2024qwen2]. Across all experiments, we set $w_{\text{max}}=0.2$ and $w_{\text{min}}=0.0$, the sampling radius of the noise term $\sigma$ to $0.01$, and the threshold in CGM to $\alpha=0.5$. We perform all training on 4 nodes, each equipped with 8 H100 GPUs. Further details are provided in Appendix C.3.
Baselines.
We compare ADHint with several baselines: (1) GRPO; (2) SFT, trained for two epochs on the full training set; (3) SFT $\rightarrow$ RL, which applies two epochs of SFT followed by GRPO; (4) LUFFY [yan2025learning], which performs on-policy and off-policy updates jointly at each step; (5) StepHint [zhang2025stephint], which mixes hint-rollouts with multiple hint ratios, naive-rollouts, and ground-truth trajectories into a group for update; (6) HintGRPO [Huang_2025_ICCV], which progressively increases the hint ratio for the hardest samples to obtain positive trajectories; and (7) GHPO [liu2025ghpo], which treats hints as part of the query and uses multiple hint ratios for the hardest samples. We note that several baselines crash during training, and for these methods, we report results from the last checkpoint available before the crash. More detailed information about the baselines and their reimplementation is provided in Appendix C.4 and Appendix C.5, respectively.
Bench. Metric GRPO HintGRPO ADHint $\Delta$
MathVista p@1 74.6 64.9 74.0 0.6
a@8 78.8 64.4 79.6 0.8
MathVerse p@1 63.1 64.2 75.2 12.1
a@8 79.0 68.7 80.1 1.1
We-Math p@1 73.1 68.6 76.4 3.3
a@8 82.6 71.7 85.6 3.0
MMMU p@1 61.0 47.2 64.4 3.4
a@8 71.7 55.5 70.1 1.6
MMMU-Pro p@1 46.5 38.4 51.8 5.3
a@8 57.9 46.0 57.9 0.0
LogicVista p@1 49.4 43.1 56.3 6.9
a@8 63.0 51.9 64.9 1.9
Avg. p@1 61.3 54.4 66.4 5.1
a@8 72.2 59.7 73.0 0.8
Table 2: Performance comparison of ADHint and baselines on Qwen3-VL-8B-Instruct. We report p@1 (pass@1) and a@8 (avg@8) for each method. The last column shows the absolute performance differences of ADHint relative to GRPO, where green indicates improvement and red indicates degradation.
Figure 6: Performance comparison of ADHint and GRPO on Qwen2.5-Math-7B. For AIME and AMC, we report avg@32 due to their small sizes, and we report pass@1 for the remaining benchmarks.
Method MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista Avg.
[…]
w/o AH-SDP 74.1/73.5 58.7/58.9 68.6/67.4 54.6/55.0 41.3/40.3 48.2/48.4 57.6/57.3
w/o AE-RDP 74.3/74.1 61.5/60.0 67.5/68.4 55.9/56.1 41.2/41.1 45.8/46.3 57.7/57.7
w/o CGM 73.1/72.8 53.9/55.0 68.2/68.0 56.3/51.1 40.0/40.0 50.5/48.2 57.0/55.9
w/o Selective Masking 74.7/73.7 58.5/61.0 67.4/67.6 55.8/55.9 41.6/40.9 47.3/46.3 57.6/57.6
Table 3: Ablation study of proposed ADHint. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
Setting MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista Avg.
$w_{\text{max}}=0$ 73.4/73.4 55.3/59.2 69.8/68.8 55.3/55.6 40.4/40.0 44.9/46.4 56.5/57.2
[…]
$w_{\text{max}}=0.4$ 73.8/73.5 62.6/57.9 68.2/67.5 54.8/54.5 41.3/40.1 47.3/47.5 58.0/56.8
$w_{\text{max}}=0.6$ 73.8/72.4 56.9/59.7 65.5/67.7 54.9/55.5 39.3/40.2 46.9/47.5 56.2/57.2
Table 4: Effect of the maximum hint ratio $w_{\text{max}}$. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
Setting MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista Avg.
$\alpha=0.25$ 72.0/72.8 56.0/59.5 67.9/68.6 55.3/55.6 40.9/40.7 48.9/48.1 56.8/57.6
[…]
$\alpha=0.75$ 76.0/75.1 53.4/62.8 66.3/68.6 55.2/56.3 40.3/41.2 47.5/47.4 56.5/58.6
Table 5: Effect of the threshold $\alpha$ in CGM. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
5.2 Main Results
SOTA performance with Qwen2.5-VL.
In Table 5, we compare the proposed ADHint against several baselines on Qwen2.5-VL-7B and Qwen2.5-VL-3B. ADHint delivers consistent performance gains at both model scales. At the 7B scale, ADHint improves pass@1 by 2.1% and avg@8 by 2.0% over GRPO. Relative to recent methods, it delivers larger gains of 6.8% in pass@1 and 11.1% in avg@8. ADHint also performs well at the 3B scale, improving pass@1 by 2.3% and avg@8 by 2.1% over GRPO, and by 9.7% and 10.5% over recent methods, respectively. These results confirm ADHint’s advanced multimodal reasoning ability and suggest that integrating difficulty into both the hint-ratio schedule and relative-advantage estimation provides a feasible solution for hint-based on-policy learning on challenging multimodal tasks. In Appendix D, we provide a more in-depth comparative analysis of the performance differences between ADHint and the baselines.
Performance with Stronger Backbone.
Table 2 presents a comparison between ADHint and several baselines on Qwen3-VL-8B-Instruct, a recent strong backbone that achieves state-of-the-art performance among models of similar scale across multiple multimodal reasoning benchmarks. Compared with GRPO, ADHint achieves competitive performance on most metrics, with absolute gains of 5.1% in pass@1 and 0.8% in avg@8. Notably, ADHint achieves gains of over 5.0% in pass@1 on MathVerse, MMMU-Pro, and LogicVista, which cover challenging problems from three different domains. These results highlight the considerable promise of ADHint for improving general multimodal capabilities. As future work, we will collect more challenging datasets tailored to the Qwen3-VL series and leverage hints from stronger models to further unlock ADHint’s potential.
Performance with LLM.
In Figure 6, we further evaluate the effectiveness of ADHint on text-only reasoning. ADHint improves the average accuracy by 2.3% over GRPO across the six math reasoning benchmarks. On OlympiadBench, which covers Olympiad-level questions, and AIME25, which consists of challenging problems from the 2025 AIME, ADHint surpasses GRPO by 5.7% and 2.4%, respectively. These results demonstrate the strong generalization of ADHint, which broadens the capability boundaries of models across different modalities.
5.3 Training Dynamics
Figure 5 presents the training dynamics of several key metrics. Figure 5(a) shows that ADHint consistently maintains rewards within a moderate regime, yielding low-variance update signals. Figures 5(b) and 5(c) indicate that ADHint maintains exploration within a healthy entropy range with more stable gradient updates. Figures 5(d) and 5(e) show that ADHint steadily increases response length, while the proportion of overlength responses does not increase, which usually indicates repetition or garbled outputs. These trends suggest that the policy model continuously acquires effective long-term thinking capabilities. Taken together, these observations indicate that ADHint stably and adaptively learns reasoning ability from off-policy hints.
5.4 Ablation Study
Table 3 reports the ablation results of four critical components in ADHint. Specifically, “w/o AH-SDP” replaces our adaptive hint-ratio schedule with an annealing hint-ratio strategy, while “w/o AE-RDP” pools hint-rollouts and naive-rollouts into a single group for advantage estimation. Removing any component results in performance degradation on both pass@1 and avg@8, underscoring the importance of our design. The difficulty-aware design in hint-ratio schedule and relative-advantage estimation ensures efficient learning that balances exploration and imitation. The gradient modulation mechanisms prevent biased and destructive updates, ensuring healthy policy learning.
5.5 Further Analysis
20% hint is sufficient to guide reasoning.
In Table 4, we vary $w_{\text{max}}$ and observe that performance is optimal when it is set to 0.2. Both increasing and decreasing the hint ratio lead to performance degradation. As shown in Figure 5(a), a 20% hint ratio is sufficient to keep rewards within the range of 0.5–0.7, providing stable learning signals at a moderate difficulty level. Further increasing the hint ratio yields excessively high rewards for hint-rollouts, which introduces larger update bias. These results suggest that the prefix segment of reasoning trajectories is crucial for guiding the policy model toward effective reasoning.
Hint-token gradients should be restricted to a reasonable range.
In Table 5, we investigate the CGM threshold $\alpha$, which controls the gradient contributions of hint tokens. The results indicate that setting $\alpha$ to 0.5 yields the best overall performance. Allowing hint tokens with excessively high or low entropy to participate in gradient updates causes the policy to deviate markedly from its initial distribution. In contrast, an overly small threshold limits effective learning from off-policy hints and reduces the efficiency of knowledge expansion.
6 Conclusion
In this work, we introduce ADHint, which explicitly incorporates difficulty into hint-based on-policy learning. Enhanced by critical components that schedule the hint ratio with respect to the sample difficulty prior, estimate relative advantages with respect to the rollout difficulty posterior, and modulate token-level gradients based on distribution consistency, ADHint surpasses existing methods of comparable scale across various benchmarks, showcasing advanced reasoning abilities and out-of-distribution generalization. Despite its impressive performance, ADHint’s scaling behavior on larger models with 32B or more parameters, as well as its generalization to tasks without verifiable answers, merits further investigation.
[…]
Supplementary Material
Our supplementary material is organized as follows:
  • Section A presents the detailed workflow of ADHint.
  • Section B provides a detailed discussion of related work.
  • Section C provides additional details on training data construction, implementation details, and the reimplementation of baselines.
  • Section D offers an in-depth comparison between ADHint and the baselines, analyzing the sources of their performance differences.
  • Section E presents several case studies that provide a more intuitive view of ADHint’s behavior.
A Training Algorithm of ADHint
In Algorithm 1, we provide the detailed workflow of ADHint as a complement to the main text.
Algorithm 1 ADHint: ADaptive Hints with Difficulty Priors for Reinforcement Learning
0:  Initial policy model $\pi_{\theta}$, training set $\mathcal{D}$, hyperparameters $w_{\text{max}},w_{\text{min}},\alpha$
0:  Optimized policy model $\pi_{\theta}^{*}$
  for $\text{step}=1,\ldots,S$ do
   Sample a mini-batch $\mathcal{D}_{b}\subset\mathcal{D}$
   for each $q\in\mathcal{D}_{b}$ do
    Generate naive-rollouts $o_{1:n}\sim\pi_{\theta}(\cdot\mid q)$
    Compute $\text{Diff}_{\text{N}}$ by Eq. 1
    Compute hint ratio $w$ by Eq. 2
    Generate hint-rollouts $o_{n+1:n+m}\sim\pi_{\theta}(\cdot\mid q,o_{i,<h})$
    Compute $\text{Diff}_{\text{H}}$ by Eq. 6
    Compute relative advantages $\tilde{A}_{i}$ by Eq. 7
    for each hint-rollout $o_{i}\in\{o_{n+1},\ldots,o_{n+m}\}$ do
     Compute token entropy by Eq. 3
     Compute continuation entropy by Eq. 4
     Compute token-wise factors $k_{i,t}$ by Eq. 5
    end for
   end for
   Update $\theta$ by ascending $\mathcal{J}_{\text{ADHint}}^{*}(\pi_{\theta})$ in Eq. 8
  end for
  return $\pi_{\theta}^{*}$
B Detailed Related Work
Multimodal Reinforcement Learning
Group Relative Policy Optimization (GRPO) [guo2025deepseek, shao2024deepseekmath] removes the critic network from PPO [schulman2017proximal] and estimates advantages through intra-group sampling. This approach has become a cornerstone for the development of recent reasoning models [yang2025qwen3, zeng2025glm, team2025kimi, team2025longcat, ye2025limo, wang2025reinforcement, tan2025towards]. In parallel, the multimodal community has explored various strategies to enhance the reasoning ability of MLLMs. Some works refine the reasoning style by encouraging models to “think with images” [zheng2025deepeyes, su2025pixel], explicitly grounding reasoning in visual features and directing attention to specific objects or regions. VL-ReThinker [wang2025vl] introduces an explicit reflection phase into the reasoning process, allowing the model to revisit and revise its initial thoughts. Other methods inject noise into rollouts [liu2025noisyrollout, wang2025perception] to improve training robustness and mitigate hallucinations. There are also approaches that design more precise reward functions for tasks such as object detection and semantic segmentation [liu2025visionreasoner, liu2025visual], as well as methods that select or curate training data to improve efficiency [li2025truth, liang2025modomodomultidomaindatamixtures]. However, these advances essentially remain optimizations of on-policy RL, primarily unlocking the latent potential of pretrained models, and they often fail to expand the models’ fundamental capability boundary.
Reinforcement Learning with Off-Policy Data
Recent work has investigated how to incorporate off-policy data into RL-based post-training. Some approaches directly integrate off-policy data into the SFT or RL objective [yan2025learning, ma2025learning, dong2025rl, fu2025srft], but they struggle with the substantial distribution mismatch between off-policy data and the current policy model. To alleviate this issue, some methods design corrections for importance sampling [yan2025learning, dong2025rl], control the trade-off between SFT and RL losses via entropy-based schedules [fu2025srft], or alternate between RL and SFT phases [ma2025learning]. Hint-based methods address this problem more flexibly by guiding the model’s reasoning with partial hints extracted from off-policy data while preserving its exploration behavior. UFT [liu2025uft] and Prefix-RFT [huang2025blending] employ the same temporally annealed hint ratios for all samples. Other methods reserve hints only for the hardest samples and determine the hint ratio from a fixed, predefined set via uniform sampling [chen2025data], binary search [zhang2025bread], or progressive search [Huang_2025_ICCV, liu2025ghpo]. StepHint [zhang2025stephint] samples rollouts under multiple hint ratios and merges the resulting hint-rollouts, naive-rollouts, and ground truth off-policy data into a single group for advantage estimation. However, these methods overlook difficulty as a key factor, leading to insufficient and unstable training. In this work, we propose ADHint, which explicitly incorporates the sample difficulty prior into the hint-ratio schedule and the rollout difficulty posterior into relative-advantage estimation. This difficulty-aware design stabilizes training, enhances exploration, and expands the model’s knowledge boundary when learning from off-policy hints.
(a) Avg@8 distribution of Qwen2.5-VL-7B (base model) on the 34K filtered samples.
(b) Distribution of samples for which Qwen3-VL-30B-A3B-Thinking fails to produce any correct hint among eight candidates.
(c) Avg@8 distribution of Qwen3-VL-235B-A22B-Thinking on the remaining hard samples.
(d) Hint-length distribution and base model accuracy after hint sampling.
Figure 7: Accuracy and length statistics during ViRL-Hint17k construction. At each sampling stage, we independently generate eight reasoning trajectories with temperature 1.0.
C Detailed Experimental Settings
C.1 Training data
For MLLM training, we construct an off-policy hint dataset based on ViRL39K [wang2025vl], which covers STEM, chart reasoning, document reasoning, and spatial reasoning. We first remove samples that involve multi-image reasoning, text-only reasoning, or format errors, obtaining roughly 34K examples. Figure 7(a) shows the avg@8 distribution of Qwen2.5-VL-7B on this filtered set. We then sample hints with Qwen3-VL-30B-A3B-Thinking, and Figure 7(b) reports the distribution of examples for which this model fails to produce any correct hint among eight candidates. For these hard cases, we further sample hints with Qwen3-VL-235B-A22B-Thinking, whose avg@8 distribution is shown in Figure 7(c). Among samples with at least one correct hint, we discard those with hint length exceeding 8K tokens and randomly select one of the eight candidates as the final hint, with the resulting hint-length distribution shown in Figure 7(d). Finally, we remove samples that are excessively simple for Qwen2.5-VL-7B and obtain ViRL-Hint17k, which we will open-source. For LLM training, we sample 21K examples with a maximum hint length of 4K tokens from public datasets [yan2025learning].
C.2 Benchmarks
For MLLM evaluation, we employ six benchmarks that cover mathematical reasoning, multidisciplinary multimodal reasoning, and logical reasoning:
  • MathVista — A consolidated benchmark for mathematical reasoning in visual contexts, which unifies math-focused and vision–language tasks into a single suite and requires fine-grained visual understanding and compositional reasoning. We use the testmini split for evaluation.
  • MathVerse — A holistic visual math benchmark that is crafted to test whether multimodal models truly understand mathematical diagrams, providing multiple variants of each problem with different modality configurations to probe robust diagram understanding. We use the Vision-Only subset.
  • We-Math — A large-scale benchmark that investigates human-like mathematical reasoning in MLLMs, decomposing problems into sub-problems and introducing multi-dimensional metrics to analyze knowledge acquisition, generalization, and rote memorization behaviours. We use its testmini split.
  • MMMU — A massive multi-discipline multimodal understanding benchmark that spans college-level subjects such as science, engineering, humanities, and business, and demands deliberate integration of visual and textual information for complex problem solving. We use the validation split.
  • MMMU-Pro — A more challenging variant of MMMU that filters out text-only questions, augments candidate options, and introduces vision-only settings, thereby providing a more robust assessment of genuine multimodal understanding and reasoning. We use the standard subset with 10 options and the vision subset to compute the overall score.
  • LogicVista — A benchmark that targets integrated logical reasoning in visual environments, focusing on compositional, step-by-step reasoning over diagrams and structured visual scenes.
For LLM evaluation, we employ six mathematical reasoning benchmarks that align with our training domain:
  • AIME24 — A dataset that contains 30 competition-style problems from the 2024 American Invitational Mathematics Examination (AIME), a well-known high-school contest featuring challenging short-answer problems.
  • AIME25 — The 2025 counterpart of AIME24, consisting of 30 similarly challenging problems from the 2025 AIME, which serves as an independent but distribution-matched test set.
  • AMC — A benchmark constructed from the American Mathematics Competitions, covering a broad range of high-school mathematics topics and providing multiple-choice questions that emphasize conceptual understanding and problem-solving skills.
  • MATH500 — A high-difficulty math competition benchmark containing 500 problems across multiple subjects, designed to assess advanced problem-solving and multi-step symbolic reasoning.
  • Minerva — A quantitative reasoning benchmark with around 500 carefully curated math problems that require multi-step derivations and the integration of textual descriptions with formal mathematical manipulations.
  • OlympiadBench — An Olympiad-level benchmark that comprises thousands of mathematically and scientifically challenging problems, offering a rigorous testbed for models’ higher-order reasoning and abstraction abilities.
C.3 Implementation Details
As a supplement to the main text, Tables 6 and 7 summarize the training configs of the GRPO and SFT baselines, respectively, while Table 8 reports the additional configs of ADHint. Note that both GRPO and ADHint use 16 rollouts to ensure a fair comparison.
C.4 Baselines
We compare ADHint against a diverse set of baselines to validate its effectiveness. These include standard on-policy RL, supervised fine-tuning, and recent methods that attempt to incorporate off-policy data into on-policy RL training:
  • GRPO — an on-policy RL method that estimates group-level relative advantages from rollouts and relies solely on an outcome reward function or model.
  • SFT — a supervised baseline that directly trains the model to imitate complete off-policy trajectories (full hints in our setting), which encourages memorization of both knowledge and response style.
  • SFT $\rightarrow$ RL — a common cold-start strategy that first performs SFT and then applies RL, serving as a strong baseline that combines the benefits of SFT and RL.
  • LUFFY — a method that augments GRPO by incorporating off-policy reasoning trajectories into the GRPO objective.
  • StepHint — a hint-based method that mixes hint-rollouts under multiple hint ratios, naive-rollouts, and off-policy reasoning trajectories into a single GRPO group for advantage estimation.
  • HintGRPO — a method that focuses on the hardest samples where the current policy fails to obtain a positive rollout. It progressively increases the hint ratio ($1/3$, $2/3$) once a positive rollout is found, followed by a one-step policy update.
  • GHPO — a method that treats hints as part of the query and excludes hint tokens from loss computation. It similarly increases the hint ratio (25%, 50%, 75%) once a positive rollout is found.
For methods that do not support MLLM training or whose code is not publicly available, we carefully reproduce them within the EasyR1 framework [zheng2025easyr1]. We note that several baselines exhibit instability and crash during training. We provide a detailed analysis in Section D.
Model Config Value
Qwen2.5-VL & Qwen3-VL max prompt length 4096
max response length 8192
temperature 1.0
learning rate 1e-6
global batchsize 128
micro batchsize 64
num rollouts 16
epoch 2
min pixels 2562828256*28*28
max pixels 128028281280*28*28
Qwen2.5-Math max prompt length 1024
max response length 4096
temperature 1.0
learning rate 1e-6
global batchsize 128
micro batchsize 64
num rollouts 16
epoch 1
Table 6: Training configs for GRPO baselines.
Method Config Value
SFT learning rate 5e-5
lr schedule cosine
warmup ratio 0.1
batchsize 64
epoch 2
min pixels 2562828256*28*28
max pixels 128028281280*28*28
Table 7: Training configs for SFT baseline.
Method Config Value
ADHint warmup steps 5
num of naive-rollouts $n$ 8
num of hint-rollouts $m$ 8
minimum hint ratio $w_{\text{min}}$ 0.0
maximum hint ratio $w_{\text{max}}$ 0.2
sampling radius of the noise term $\theta$ 0.01
threshold of CGM $\alpha$ 0.5
Table 8: Additional training configs for ADHint.
C.5 Reimplementation Details of Baselines
For all baselines, we strictly follow the official implementations and the settings reported in the original papers. We train each method for 2 epochs on the same ViRL-Hint17k dataset to ensure a fair comparison. Below, we describe the reimplementation details for each baseline.
LUFFY.
Following the paper, we add an off-policy objective to the standard GRPO objective to obtain the mixed-policy training. We also reimplement the policy shaping module to modulate the update gradients of off-policy tokens and set the shaping hyperparameter to the default value of 0.1 reported in LUFFY. We omit both the KL divergence constraint and the standard advantage normalization.
StepHint.
We obtain hint-rollouts under hint ratios of 25%, 50%, and 75%, and group them with naive-rollouts and complete off-policy reasoning trajectories into a single GRPO group for relative advantage estimation. We omit the KL constraint and mask hint tokens during optimization for negative hint-rollouts.
HintGRPO.
We set the group number to $M=3$, so the hint ratio is progressively increased from 0 to $1/3$ and then $2/3$ until a positive rollout appears. We set the KL coefficient to 0.04 and the maximum response length to 8K tokens. This setting matches our ViRL-Hint17k dataset, which has a relatively long average hint length (around 3100 tokens) and is substantially longer than that in the original HintGRPO experiments. We do not reproduce the Text-Bias Calibration module from HintGRPO because it is a test-time method for correcting perception bias and mitigating hallucinations rather than a training-time hint-based component.
GHPO.
For the hardest samples where the current policy fails to produce a correct answer, we insert hints into the user prompt using the official prompt template. The hint ratio is progressively increased from 25% to 50% and then to 75% until a positive rollout appears. We then use the corresponding hint-rollouts as the update signal for that sample. Following the original paper, we omit the KL constraint and apply a 5-step GRPO warmup.
D Comparative Analysis of ADHint and Baselines
As shown in Table 5 and Figure 2, ADHint consistently outperforms all baselines. Here, we provide a detailed comparative analysis of ADHint and these baselines to better understand the sources of their performance differences.
SFT learns, RL generalizes.
Compared with pure GRPO, the $\mathrm{SFT}\rightarrow\mathrm{RL}$ pipeline improves avg@8 by 1.3% but reduces pass@1 by 12.2% at the 7B scale. This pattern indicates that SFT injects new knowledge into the model and raises the upper bound of its performance when multiple samples are allowed, whereas the subsequent RL phase, starting from an SFT-initialized policy, tends to lose part of the exploration capability that is crucial for strong pass@1 performance compared with directly training with on-policy RL. In practice, SFT encourages the model to follow a specific response style, which makes the response length and entropy during RL training concentrate in a relatively narrow range, thereby limiting exploration across diverse reasoning trajectories. In contrast, ADHint achieves the best results on both pass@1 and avg@8. By explicitly accounting for difficulty, ADHint assigns an appropriate hint ratio to each sample and estimates fair relative advantages for each rollout, balancing exploration and imitation. As a result, the policy model simultaneously strengthens its reasoning ability and absorbs external knowledge from off-policy hints.
Training collapses when directly using full off-policy data.
During our reimplementation of LUFFY [yan2025learning], we observe that training collapses at an early stage (around 40 steps), accompanied by a sharp increase in both entropy and response length. We attribute this to over-imitation of the off-policy distribution, especially when there is a large discrepancy between the model that generates the off-policy data and the current policy. LUFFY feeds the complete off-policy trajectories into the mixed-policy GRPO objective as the off-policy term, while the other term remains the original on-policy GRPO objective. This forces the policy model to simultaneously fit two highly mismatched distributions, making training unstable and prone to collapse. In contrast, ADHint treats difficulty as a vital signal. It adjusts the hint ratio based on sample difficulty and estimates relative advantages based on rollout difficulty, achieving a better trade-off between learning from off-policy hints and maintaining exploratory on-policy reasoning.
Fixed hint ratios always lead to poor generalization.
GHPO [liu2025ghpo] and HintGRPO [Huang_2025_ICCV] share a similar design philosophy: they add hints with fixed ratios (e.g., 25%, 50%) only to the hardest problems. The key difference is that GHPO injects hints into the prompt and excludes them from the training loss, whereas HintGRPO treats hints as part of the response and includes them in the loss computation. In our experiments, we observe that GHPO exhibits large update magnitudes and highly unstable training, likely because the model is not truly learning the knowledge contained in the hints but instead adapting to an out-of-distribution reasoning pattern. In contrast, HintGRPO trains more stably and maintains a high average reward during training. However, as shown in Table 5, HintGRPO performs poorly on the benchmarks and shows limited generalization. The progressive hinting strategy for the hardest samples helps obtain positive rollouts, but it ignores sample difficulty and can provide excessive guidance, causing SFT-like overfitting to the hint distribution and weakening exploration. As a result, the model attains inflated training rewards under heavy hinting but fails to translate these gains into improvements on the benchmarks. By comparison, ADHint leverages a sample difficulty prior to assigning appropriate guidance to each example and uses low-variance update signals to gradually push the model’s capability boundary, enabling it to learn from off-policy data while preserving its inherent exploration.
Imbalanced intra-group advantage estimation leads to abnormal updates.
StepHint [zhang2025stephint] mixes hint-rollouts with multiple fixed hint ratios, naive-rollouts, and ground-truth trajectories into a single group for advantage estimation without distinguishing among them. As shown in Figure 1(b) of the main text, this imbalanced intra-group estimation allows longer hint-rollouts with more positive trajectories to receive disproportionately large advantages, which biases policy updates toward imitating hints and leads to abnormal update behavior. This issue becomes more severe on ViRL-Hint17k, where the average hint length is about 3,100 tokens, while Qwen2.5-VL models typically generate only 400–500 tokens. In contrast, ADHint introduces AE-RDP, which estimates rollout advantages using the difficulty posterior evaluated from naive-rollouts and hint-rollouts, balancing learning between the two rollout types. CGM and Selective Masking further modulate token-level gradients within hints, which promotes more stable and reliable policy updates.
E Case Study
Figure 8 shows a representative training example. This case includes the extra prompt we add beyond the user query, which instructs the model to think before producing its response. For clarity, we omit this additional prompt in the subsequent cases. We see that the naive-rollout of the policy model fails to produce the correct answer. However, when we give the policy a hint taken from the complete off-policy reasoning trajectory, it continues the reasoning process and computes the final answer $z=2\sqrt{15}$. In this way, the model not only learns the essential knowledge required to solve the problem from the hint, but also improves its reasoning ability.
Figure 9 shows two representative benchmark examples. In the top case, the GRPO model lacks sufficient geometric knowledge and reaches an incorrect conclusion that “a circle can tile a plane because it forms a complete 360-degree angle when placed around a point.” In contrast, the ADHint model performs careful reasoning: it first interprets the meaning of “tile a plane” in the question, and then, through step-by-step reasoning, reaches the correct conclusion that “the only shape that can tile a plane is the right triangle.” In the bottom case, the GRPO model produces flawed reasoning, such as claiming that “E is directly above B,” and ultimately predicts the wrong option “C”. By comparison, the ADHint model conducts logically coherent reasoning: it first “visualizes the solid shape” and then gradually rules out incorrect options, finally selecting the correct answer “D”. These case studies indicate that GRPO fails to acquire the necessary external knowledge and exhibits weak reasoning ability, whereas ADHint treats difficulty as a key factor, learns essential knowledge from hints in a controllable manner, and develops more structured logical reasoning. This demonstrates ADHint’s strong ability to effectively integrate off-policy data into on-policy RL training.
Figure 8: A training example comparing naive-rollouts and hint-rollouts in ADHint.
Figure 9: Two benchmark examples comparing ADHint with the GRPO baseline.

### v2

@@
-ADHint: Adaptive Hints with Difficulty Priors for Reinforcement Learning
+
 Abstract
-To combine the advantages of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL), recent methods have integrated “hints” into post-training, which are prefix segments of complete reasoning trajectories, aiming for powerful knowledge expansion and reasoning generalization.
-However, existing hint-based RL methods typically ignore difficulty when scheduling hint ratios and estimating relative advantages, leading to unstable learning and excessive imitation of off-policy hints.
-In this work, we propose ADHint, which treats difficulty as a key factor in both hint-ratio schedule and relative-advantage estimation to achieve a better trade-off between exploration and imitation.
-Specifically, we propose Adaptive Hint with Sample Difficulty Prior, which evaluates each sample’s difficulty under the policy model and accordingly schedules an appropriate hint ratio to guide its rollouts.
-We also introduce Consistency-based Gradient Modulation and Selective Masking for Hint Preservation to modulate token-level gradients within hints, preventing biased and destructive updates.
-Additionally, we propose Advantage Estimation with Rollout Difficulty Posterior, which leverages the relative difficulty of rollouts with and without hints to estimate their respective advantages, thereby achieving more balanced updates.
-Extensive experiments across diverse modalities, model scales, and domains demonstrate that ADHint delivers superior reasoning ability and out-of-distribution generalization, consistently surpassing existing methods in both pass@1 and avg@8.
-Our code and dataset will be made publicly available upon paper acceptance.
-*: Equal contribution, 🖂: Corresponding author.
+To address the limited capability expansion and low sample efficiency of Reinforcement Learning (RL), recent methods have integrated “hints” into post-training, which are prefix segments of complete reasoning trajectories, aiming for powerful knowledge expansion and reasoning generalization.
+However, existing hint-based RL methods often neglect the role of difficulty in the hint-ratio schedule and relative-advantage estimation, resulting in unstable learning and excessive imitation of off-policy hints.
+To address this, we propose ADHint, which explicitly integrates difficulty into both processes to achieve a better trade-off between exploration and imitation.
+Specifically, we propose Adaptive Hint with Sample Difficulty Prior, which evaluates the difficulty of each sample under the current policy to schedule an appropriate hint ratio for rollout generation.
+Furthermore, we introduce Consistency-based Gradient Modulation alongside Selective Masking for Hint Preservation, which jointly modulate token-level gradients within hints to prevent biased and destructive updates.
+Additionally, we propose Advantage Estimation with Rollout Difficulty Posterior, which leverages the relative difficulty of rollouts with and without hints to compute their respective advantages, yielding more balanced updates.
+Extensive experiments across diverse modalities, model scales, model families, and domains demonstrate that ADHint achieves superior reasoning capabilities and out-of-distribution generalization.
+Code and datasets will be made publicly available upon paper acceptance.
 1 Introduction
@@
 Recent advancements in Large Language Models (LLMs) [yang2024qwen2, yang2025qwen3, li2025system, zeng2025glm] and Multimodal Large Language Models (MLLMs) [bai2025qwen3vltechnicalreport, liu2025visual, huang2025vision, wang2025internvl3] highlight the potential to enhance reasoning capabilities by large-scale Reinforcement Learning with Verifiable Rewards (RLVR) [shao2024deepseekmath, guo2025deepseek, yu2025dapo], such as Proximal Policy Optimization (PPO) [schulman2017proximal] and Group Relative Policy Optimization (GRPO) [shao2024deepseekmath, guo2025deepseek].
-However, RLVR primarily reinforces existing behaviors rather than instilling novel reasoning abilities [chusft, li2025system].
-On the contrary, Supervised Fine-Tuning (SFT) offers a complementary approach to facilitating the incorporation of new knowledge and reasoning patterns [chen2025sft, yan2025learning].
-Nevertheless, the SFT models often exhibit limited generalization to out-of-distribution scenarios.
-These contrasting characteristics of RL and SFT motivate a promising research direction: How can RL and SFT effectively collaborate to enhance both reasoning abilities and out-of-distribution generalization?
-To achieve this goal, early methods [yan2025learning, ma2025learning, dong2025rl, fu2025srft] attempt to incorporate on-policy RL with full reasoning trajectories to internalize external knowledge.
-However, these methods often struggle with insufficient exploration [wu2025templaterlstructuredtemplateguidedreinforcement, xi2025bapostabilizingoffpolicyreinforcement] and even training collapse [liang2025squeezesoakedspongeefficient, cui2025entropymechanismreinforcementlearning].
-Recent methods [Huang_2025_ICCV, liu2025uft, zhang2025stephint] introduce hints, which refer to the prefix segments of complete reasoning trajectories, allowing the policy model to explore the continuation rather than reinforcing the model to mimic the entire sequence.
-Specifically, these methods assign fixed [zhang2025stephint] or time-varying [liu2025uft, huang2025blending] hint ratios to all samples (or only to the hardest samples [Huang_2025_ICCV, liu2025ghpo]), and then ultize hint-rollouts derived from hint-guided reasoning together with naive-rollouts generated without hints for group relative advantage estimation [zhang2025stephint, huang2025blending].
+However, current on-policy RLVR faces two major challenges: (1) Limited capability expansion: RLVR is inherently bounded by the base model.
+It primarily amplifies existing behaviors and refines known reasoning chains, rather than instilling genuinely novel reasoning abilities beyond its initial capability boundaries [chusft, li2025system].
+(2) Low sample efficiency: The learning process is bottlenecked by the current policy’s performance, yielding critically sparse reward signals that render hard samples difficult to exploit [wu2025templaterlstructuredtemplateguidedreinforcement, xi2025bapostabilizingoffpolicyreinforcement].
+To mitigate these limitations, recent methods [Huang_2025_ICCV, liu2025uft] introduce hints, defined as prefix segments of complete reasoning trajectories, which guide the policy to explore the continuation, thereby acquiring novel capabilities and solving challenging problems.
+Specifically, these methods assign fixed [zhang2025stephint] or time-varying [liu2025uft, huang2025blending] hint ratios to all samples (or only to the hardest samples [Huang_2025_ICCV, liu2025ghpo]), and then utilize hint-rollouts derived from hint-guided reasoning together with naive-rollouts generated without hints for group relative advantage estimation [zhang2025stephint, huang2025blending].
 Despite their empirical successes, existing hint-based RL methods are sensitive to sample-difficulty in the hint-ratio schedule and rollout-difficulty in the relative-advantage estimation.
 Specifically, most previous methods apply a uniform hint ratio to both easy and hard samples for model reasoning, and then generate hint-rollouts with various difficulties.
-Policy learning from such heterogeneous rollouts could introduce high-variance model optimization, and thus undermine training stability [bengio2009curriculum, wang2025dump, parashar2025curriculum].
-As shown in Figure 1(a), the model with annealing hint ratio (denoted as Baseline) meets with training collapse since its entropy rises sharply at the end of training, indicating insufficient absorption of external knowledge and a drift in the model’s own policy.
-In contrast, when introducing the hint-ratio schedule with sample-difficulty (our method, discussed later), the entropy converges stably so that the model could learn reasoning ability, not superficial patterns in off-policy hints.
+Policy learning from such heterogeneous rollouts could introduce high-variance model optimization, and thus undermine training stability [bengio2009curriculum, parashar2025curriculum, li2025adacurl].
+As shown in Figure 1(a), the model with annealing hint ratio (denoted as Baseline) suffers from training collapse since its entropy rises sharply at the end of training, indicating insufficient absorption of external knowledge and a drift in the model’s own policy.
+In contrast, when incorporating the hint-ratio schedule with sample-difficulty (our method, discussed later), the entropy converges stably, ensuring that the model internalizes robust reasoning abilities rather than superficial patterns within off-policy hints.
 As for relative-advantage estimation, most existing approaches pool hint-rollouts and naive-rollouts into a single group so that the relative-advantage is dominated by hint-rollouts, with the model learning to directly imitate the off-policy hint distribution rather than exploring with its own policy under hint guidance.
 Specifically, externally guided hint-rollouts are typically simpler than naive-rollouts, leading to more trajectories with positive advantages that dominate the update signal and bias the policy toward fitting the off-policy distribution.
-Additionally, hints from external reasoning models are usually much longer than the policy’s own outputs, further exacerbating the imbalance.
+Additionally, hints from external reasoning models are usually much longer than the policy’s own outputs, further exacerbating this imbalance.
 As shown in Figure 1(b), the average reward of hint-rollouts increases steadily, whereas that of naive-rollouts collapses, causing the model (denoted as Baseline) to eventually lose the ability to perform inference without hints.
-Figure 2: Overall comparison between ADHint and baselines.
-Each axis corresponds to a dataset–metric pair, where p@1 denotes pass@1, and a@8 denotes avg@8.
-ADHint achieves the best or highly competitive performance on most metrics.
+(a) Performance on Qwen2.5-VL-7B.
+(b) Gains across scenarios.
+Figure 2: Performance Comparison and Generalization.
+(a) Radar chart on Qwen2.5-VL-7B-Instruct.
+Each axis represents a dataset-metric pair (p@1: pass@1, a@8: avg@8), where ADHint consistently achieves the best or highly competitive performance.
+(b) Absolute gains over the best baselines.
+ADHint demonstrates robust improvements across diverse modalities, model scales, model families, and domains.
 To address these issues, we propose ADHint (ADaptive Hints with Difficulty Priors for Reinforcement Learning), which incorporates difficulty into both the hint-ratio schedule and the relative-advantage estimation to achieve a principled trade-off between exploration and imitation.
-For the hint-ratio schedule, we propose Adaptive Hint with Sample Difficulty Prior (AH-SDP), which adaptively determines the hint ratio for each sample based on the difficulty prior estimated from the average reward of its naive-rollouts, keeping the resulting hint-rollouts within a moderate difficulty range to provide stable update signals.
-Furthermore, to prevent large shifts toward the off-policy distribution that trigger entropy collapse, we propose Consistency-based Gradient Modulation (CGM), which downweights the gradients of hint tokens whose entropy deviates significantly from that of the policy-generated continuation.
-We also introduce Selective Masking for Hint Preservation (Selective Masking), which filters out erroneous update signals from hint-rollouts with negative advantages.
+For the hint-ratio schedule, we propose Adaptive Hint with Sample Difficulty Prior (AH-SDP), which adaptively determines the hint ratio for each sample based on a difficulty prior estimated from the average reward of its naive-rollouts, maintaining the resulting hint-rollouts within a moderate difficulty range to provide stable update signals.
+Furthermore, to prevent drastic shifts toward the off-policy distribution that trigger entropy collapse, we propose Consistency-based Gradient Modulation (CGM), which downweights the gradients of hint tokens whose entropy deviates significantly from that of the policy-generated continuation.
+We also introduce Selective Masking for Hint Preservation (Selective Masking), which discards erroneous update signals from hint-rollouts with negative advantages.
 For the relative-advantage estimation, we propose Advantage Estimation with Rollout Difficulty Posterior (AE-RDP), which constructs the rollout difficulty posterior from the average rewards of both naive-rollouts and hint-rollouts, and uses it to estimate advantages.
-Positive naive-rollouts with greater difficulty receive larger advantages, as they provide valuable learning signals that are better aligned with current policy, whereas negative hint-rollouts with lower difficulty are penalized more heavily.
-We conduct extensive experiments across both multimodal and text-only settings, multiple model scales, and diverse benchmarks covering mathematical, multi-disciplinary, and logical reasoning.
-As shown in Figure 2, ADHint consistently yields substantial gains in pass@1 and avg@8, demonstrating synergistic improvements in both reasoning ability and out-of-distribution generalization.
-Experiments on the recent, stronger backbone Qwen3-VL-8B-Instruct [yang2025qwen3] further validate the effectiveness of ADHint.
+Positive naive-rollouts with greater difficulty receive larger advantages, as they provide valuable learning signals better aligned with the current policy, whereas negative hint-rollouts with lower difficulty are penalized more heavily.
+We conduct extensive experiments across multiple modalities, model scales, model families, domains, and tasks.
+As shown in Figure 2, ADHint consistently yields substantial gains, demonstrating synergistic improvements in both reasoning ability and out-of-distribution generalization.
 Our main contributions are summarized as follows:
-  • We reveal that difficulty is a crucial signal for both hint-ratio schedule and relative-advantage estimation, and that neglecting it results in unstable learning and overfitting to the off-policy distribution.
-  • We propose ADHint, which explicitly exploits sample difficulty priors and rollout difficulty posteriors for hint-ratio schedule and relative-advantage estimation, leading to a better balance between exploration and imitation.
-  • Extensive experiments across diverse settings consistently demonstrate ADHint’s strong performance, delivering significant improvements in pass@1 and avg@8 on multi-domain benchmarks.
+  • We reveal that difficulty is a crucial signal for both the hint-ratio schedule and relative-advantage estimation, and that neglecting it results in unstable learning and overfitting to the off-policy distribution.
+  • We propose ADHint, which explicitly exploits sample difficulty priors and rollout difficulty posteriors for the hint-ratio schedule and relative-advantage estimation, leading to a better balance between exploration and imitation.
+  • Extensive experiments across diverse settings consistently demonstrate ADHint’s superiority, delivering robust and significant performance gains across comprehensive multi-domain benchmarks.
 2 Related Work
 2.1 Multimodal Reinforcement Learning
-GRPO [guo2025deepseek, shao2024deepseekmath] removes the critic network from PPO [schulman2017proximal] and estimates advantages via intra-group sampling, which has become a cornerstone for recent reasoning models [yang2025qwen3, zeng2025glm, team2025kimi, team2025longcat, ye2025limo, wang2025reinforcement, tan2025towards].
-In parallel, the multimodal community has explored various strategies to enhance MLLM reasoning, including refining reasoning styles [wang2025vl, zheng2025deepeyes, su2025pixel], improving rollout procedures for training robustness [liu2025noisyrollout, wang2025perception], redesigning reward functions [liu2025visionreasoner, liu2025visual], and data augmentation [liang2025modomodomultidomaindatamixtures, li2025truth].
-Despite these advances, RL algorithms primarily focus on unlocking the latent potential of pretrained models and often fail to expand their fundamental capability boundary.
-2.2 Reinforcement Learning with Off-Policy Data
-Recent work explores incorporating off-policy data into RL-based post-training.
-Methods that directly integrate off-policy data into SFT or RL objectives [yan2025learning, ma2025learning, dong2025rl, fu2025srft] often struggle with the substantial distribution mismatch between off-policy data and the on-policy model.
-Hint-based methods alleviate this by guiding the model’s reasoning with partial hints extracted from off-policy data while preserving its own exploration.
-UFT [liu2025uft] and Prefix-RFT [huang2025blending] employ temporally annealed hint ratios, while StepHint [zhang2025stephint] mixes hint-rollouts from multiple hint ratios with naive-rollouts and ground-truth into a single group.
-Other methods reserve hints only for the hardest samples and determine the hint ratio via uniform sampling [chen2025data], binary search [zhang2025bread], or progressive search [Huang_2025_ICCV, liu2025ghpo].
+GRPO [guo2025deepseek, shao2024deepseekmath] removes the critic network from PPO [schulman2017proximal] and estimates advantages via intra-group sampling, serving as a cornerstone for recent reasoning models [yang2025qwen3, zeng2025glm, team2025kimi, team2025longcat, ye2025limo, wang2025reinforcement, tan2025towards].
+In parallel, the multimodal community has explored various strategies to enhance MLLM reasoning, including refining reasoning styles [wang2025vl, zheng2025deepeyes, su2025pixel], improving rollout procedures for training robustness [liu2025noisyrollout, wang2025perception], redesigning reward functions [liu2025visionreasoner, liu2025visual], and augmenting training data [liang2025modomodomultidomaindatamixtures, li2025truth].
+Despite these advances, RL algorithms mainly unlock the latent potential of pretrained models and often fail to expand their fundamental capability boundary.
+2.2 Reinforcement Learning with Off-policy Hints
+Recent works incorporate off-policy hints into RL-based post-training by providing partial solution prefixes from off-policy data to guide the model’s reasoning while preserving on-policy exploration.
+UFT [liu2025uft] and Prefix-RFT [huang2025blending] employ temporally annealed hint ratios, whereas StepHint [zhang2025stephint] mixes hint-rollouts at multiple hint ratios with naive-rollouts and ground-truth trajectories within a single group.
+SEELE [li2025staying] introduces learnable parameters to predict instance-specific hint ratios in a multi-round training scheme.
+Other methods reserve hints exclusively for the most challenging samples and determine the hint ratio via uniform sampling [chen2025data], binary search [zhang2025bread], or progressive search [Huang_2025_ICCV, liu2025ghpo].
 In this work, we propose ADHint, which explicitly incorporates difficulty into both the hint-ratio schedule and the relative-advantage estimation to address the unstable learning and excessive imitation observed in existing hint-based methods.
-More detailed discussions are provided in Appendix B.
-3 Preliminary
-Figure 3: Overview of ADHint.
+More detailed discussions are provided in Appendix C.
+Figure 3: An overview of ADHint.
 For each sample, ADHint successively infers twice to obtain naive-rollouts and hint-rollouts, where the hint ratio of the latter is scheduled based on the difficulty evaluated from the former.
 It then leverages the relative difficulty of naive-rollouts and hint-rollouts to estimate their respective advantages, and modulates the gradient of each hint token based on the consistency of entropy.
+3 Preliminary
 3.1 Group Relative Policy Optimization
@@
 The current policy model $\pi_{\theta}$ is optimized by maximizing the following objective:
-$\mathcal{J}_{\text{GRPO}}(\pi_{\theta})={}\mathbb{E}_{q\sim Q,{\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}}$
-$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Biggl\{\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i}\;-\beta\,D_{\mathrm{KL}}\bigl(\pi_{\theta}\,\|\,\pi_{\text{ref}}\bigr)\Biggr\}.$
+$\mathcal{J}_{\text{GRPO}}(\pi_{\theta})=\mathbb{E}_{q\sim Q,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Biggl\{\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i}-\beta\,D_{\mathrm{KL}}\bigl(\pi_{\theta}\,\|\,\pi_{\text{ref}}\bigr)\Biggr\}$
 where the standard min and clip operations for avoiding extreme gradients are omitted for clarity, and $D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})$ is a KL constraint term.
@@
 The gradient for a given $q$ and $o_{i}$ is calculated as:
-$\nabla_{\theta}\mathcal{J}^{*}_{\text{GRPO}}(\pi_{\theta})\Big|_{q,o_{i}}=\frac{1}{G\cdot|o_{i}|}$
-$\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\hat{A}_{i}\cdot\nabla_{\theta}\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t})).$
+$\nabla_{\theta}\mathcal{J}^{*}_{\text{GRPO}}(\pi_{\theta})\Big|_{q,o_{i}}=\frac{1}{G\cdot|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\hat{A}_{i}\cdot\nabla_{\theta}\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t})).$
 3.2 Hint-based On-Policy Learning
@@
 Notably, since hint tokens ($t\leq h$) are not generated by the policy model, following prior work [yan2025learning], we set $\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})=1$ for all hint tokens in hint-rollouts.
-4 Method
 (a) ADHint w/o CGM
@@
 The right subfigure compares the entropy of the baseline and GRPO, where “Baseline: Hint/Naive” is computed on hint- or naive-rollouts only.
+4 Method
 4.1 Overview
 As shown in Figure 3, ADHint comprises four crucial modules: (1) Adaptive Hint with Sample Difficulty Prior (AH-SDP), where the policy model successively performs two inferences to obtain naive-rollouts and hint-rollouts, and accordingly schedules the hint ratio for the latter based on the sample difficulty prior evaluated from the former; (2) Advantage Estimation with Rollout Difficulty Posterior (AE-RDP), which uses the rollout difficulty posterior to estimate relative advantages for naive-rollouts and hint-rollouts, covering both correct and wrong outcomes; (3) Consistency-based Gradient Modulation (CGM), where the gradient of each hint token is modulated according to the consistency of its entropy distribution with that of the continuation; and (4) Selective Masking for Hint Preservation (Selective Masking), which masks the update signals of hint tokens in hint-rollouts that remain incorrect even with guidance.
-Additionally, Algorithm 1 summarizes the overall procedure of ADHint.
+Additionally, Algorithm 1 summarizes the overall procedure.
 4.2 Adaptive Hint with Sample Difficulty Prior
@@
 These observations suggest that the model tends to memorize superficial hint patterns rather than internalize the underlying knowledge.
-Unlike existing methods, we use the difficulty prior evaluated from each sample’s naive-rollouts to schedule the hint ratio, thereby keeping hint-rollouts within a moderate-difficulty regime (Figure 5(a)) and providing low-variance update signals at every step.
+Unlike existing methods, we use the difficulty prior evaluated from each sample’s naive-rollouts to schedule the hint ratio, thereby keeping hint-rollouts within a moderate-difficulty regime (Figure 4(c)) and providing low-variance update signals at every step.
 For each query $q$, we define the difficulty score of naive-rollouts as
@@
 $w\;=\;f(\text{Diff}_{\text{N}})\;=\;(w_{\text{max}}-w_{\text{min}})\cdot\text{Diff}_{\text{N}}+w_{\text{min}}+\sigma.$ (2)
-where $w_{\text{max}}$ and $w_{\text{min}}$ denote the maximum and minimum ratio hyperparameters, and $\sigma\sim\mathcal{U}(-R,R)$ is uniformly distributed noise.
+where $w_{\text{max}}$ and $w_{\text{min}}$ denote the maximum and minimum ratios, and $\sigma\sim\mathcal{U}(-R,R)$ is uniformly distributed noise.
+We fix $w_{\text{min}}=0$ since easy questions require no hints, thus eliminating it as a hyperparameter.
 Finally, the adaptively computed ratio $w$ determines the hint length $h$, which we then use to generate $m$ hint-rollouts $\{o_{n+1},\dots,o_{n+m}\}$.
@@
 Specifically, for each hint token in a hint-rollout $o_{i}$, we define the token-level entropy as
-$H_{i,t}\;=\;\mathcal{H}\!\big(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\big),$ (3)
+$H_{i,t}=\mathcal{H}(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})),$ (3)
 and the average entropy of the policy-generated continuation as
-$\bar{H}_{i}\;=\;\frac{1}{\,|o_{i,>h}|\,}\sum_{t=h+1}^{|o_{i}|}\mathcal{H}\!\big(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\big).$ (4)
+$\bar{H}_{i}=\frac{1}{|o_{i,>h}|}\sum_{t=h+1}^{|o_{i}|}\mathcal{H}(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})).$ (4)
 We then scale the gradient of each token by
-$k_{i,t}\;=\;\begin{cases}g\!\left(\dfrac{H_{i,t}}{\bar{H}_{i}}\right),&1\leq t\leq h,\\[6.0pt] 1,&t>h,\end{cases}$
+$k_{i,t}=g(H_{i,t}/\bar{H}_{i})\text{ if }1\leq t\leq h,\text{ and }1\text{ otherwise.}$
 where $g(x;\alpha)$ is a symmetric cosine-based schedule centered at $g(1)=1$, with a ramp-up on $\alpha,1]$ and a ramp-down on $1,1/\alpha]$ for $0<\alpha\leq 1$:
-$g(x;\alpha)\;=\;\begin{cases}\displaystyle\!\sin\!\Big(\tfrac{\pi}{2}\cdot\tfrac{x-\alpha}{1-\alpha}\Big),&\alpha\leq x\leq 1,\\[8.0pt] \displaystyle\!\cos\!\Big(\tfrac{\pi}{2}\cdot\tfrac{x-1}{(1/\alpha)-1}\Big),&1<x\leq 1/\alpha,\\[8.0pt] 0,&\text{otherwise}.\end{cases}$
+$g(x;\alpha)=\begin{cases}\sin\bigl(\frac{\pi}{2}\cdot\frac{x-\alpha}{1-\alpha}\bigr),&\alpha\leq x\leq 1,\\[4.0pt] \cos\bigl(\frac{\pi}{2}\cdot\frac{x-1}{1/\alpha-1}\bigr),&1<x\leq 1/\alpha,\\[4.0pt] 0,&\text{otherwise}.\end{cases}$
 Selective Masking for Hint Preservation.
@@
 Thus, we selectively mask the gradients of hint tokens in this case:
-$k_{i,t}\;=\;\begin{cases}0,&1\leq t\leq h\ \text{and}\ \hat{A}_{i}\leq 0,\\[4.0pt] g\!\left(\dfrac{H_{i,t}}{\bar{H}_{i}}\right),&1\leq t\leq h\ \text{and}\ \hat{A}_{i}>0,\\[6.0pt] 1,&t>h.\end{cases}$ (5)
+$k_{i,t}=\begin{cases}0,&1\leq t\leq h\text{ and }\hat{A}_{i}\leq 0,\\ g(H_{i,t}/\bar{H}_{i}),&1\leq t\leq h\text{ and }\hat{A}_{i}>0,\\ 1,&t>h.\end{cases}$ (5)
 To unify the notation, we fix $k_{i,t}=1$ for naive-rollouts.
@@
 In contrast, hint-rollouts are relatively easier and thus warrant stronger penalties when incorrect.
-Specifically, for each query $q$, we define the difficulty score of naive-rollouts in Eq. 1. And similarly define the difficulty score of hint-rollouts as
-$\text{Diff}_{\text{H}}\;=\;1-\mathcal{M}(n+1:)\;=\;1-\operatorname{mean}(r_{n+1},\dots,r_{n+m}).$ (6)
+Specifically, for each query $q$, we define the difficulty score of naive-rollouts in Eq. 1, and similarly define the difficulty score of hint-rollouts as
+$\mathrm{Diff}_{\mathrm{H}}=1-\mathcal{M}(n+1:)=1-\operatorname{mean}(r_{n+1},\dots,r_{n+m}).$ (6)
 Then the difficulty score of rollout $o_{i}$ is denoted as
-$\text{Diff}_{i}\;=\;\begin{cases}\text{Diff}_{\text{N}},&1\leq i\leq n,\\ \text{Diff}_{\text{H}},&n{+}1\leq i\leq n{+}m.\end{cases}$
-Let $\hat{A}_{i}$ denote the relative advantage of rollout $o_{i}$ estimated by pooling all n+mn{+}m trajectories. We define $s_{i}$ as a sign indicator:
-$s_{i}\;=\;\begin{cases}1,&\hat{A}_{i}>0,\\ -1,&\hat{A}_{i}\leq 0.\end{cases}$
+$\mathrm{Diff}_{i}=\mathrm{Diff}_{\mathrm{N}}\text{ if }1\leq i\leq n,\text{ and }\mathrm{Diff}_{\mathrm{H}}\text{ if }n+1\leq i\leq n+m.$
+Let $\hat{A}_{i}$ denote the relative advantage of rollout $o_{i}$, estimated by pooling all $n+m$ trajectories.
+We define a sign indicator $s_{i}=\operatorname{sgn}(\hat{A}_{i})$, where $s_{i}=-1$ when $\hat{A}_{i}\leq 0$.
 Finally, the relative advantages in ADHint are estimated as
-$\tilde{A}_{i}\;=\;\left(\frac{\mathcal{M}(1:n+m)}{1-\text{Diff}_{i}}\right)^{s_{i}}\hat{A}_{i}.$ (7)
+$\tilde{A}_{i}=\left(\frac{\mathcal{M}(1:n+m)}{1-\mathrm{Diff}_{i}}\right)^{s_{i}}\hat{A}_{i}.$ (7)
 4.4 Overall Gradient Computation
 Building on the above mechanisms, we compute the gradient for a given query $q$ and rollout $o_{i}$ as:
-$\nabla_{\theta}\mathcal{J}^{*}_{\text{ADHint}}(\pi_{\theta})\Big|_{q,o_{i}}=\frac{1}{G\cdot|o_{i}|}$ (8)
-$\sum_{t=1}^{|o_{i}|}k_{i,t}\cdot\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\tilde{A}_{i}\cdot\nabla_{\theta}\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t})).$
-Notably, to avoid inaccurate difficulty prior evaluation caused by formatting errors, we warm up the policy with GRPO for the first five steps so that the model follows the required response format (Figure 5(f)).
-(a) Training reward
-(b) Policy entropy
-(c) Gradient norm
-(d) Response length
-(e) Clipping ratio of long responses
-(f) Format reward
-Figure 5: Training dynamics of key metrics for ADHint and GRPO.
-“ADHint: Hint/Naive” reports metrics computed on hint- or naive-rollouts only, while “ADHint: Overall” aggregates both rollout types.
-ADHint maintains stable learning and yields concurrent improvements in reasoning generalization and knowledge expansion.
+$\nabla_{\theta}\mathcal{J}^{*}_{\text{ADHint}}(\pi_{\theta})\Big|_{q,o_{i}}=\frac{1}{G\cdot|o_{i}|}\sum_{t=1}^{|o_{i}|}k_{i,t}\cdot\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\tilde{A}_{i}\cdot\nabla_{\theta}\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t})).$
+Notably, to avoid inaccurate difficulty prior evaluations caused by formatting errors, we warm up the policy model with GRPO for the first five steps to ensure the model follows the required response format (Figure 4(h)).
 5 Experiments
+5.1 Settings
+We conduct experiments across multiple modalities, model scales, model families, and domains to validate ADHint’s effectiveness.
+For common MLLM reasoning tasks (e.g., math and chart reasoning), we evaluate ADHint on Qwen2.5-VL-7B-Instruct [bai2025qwen2], Qwen3-VL-8B-Instruct [bai2025qwen3vltechnicalreport], Qwen2.5-VL-3B-Instruct, Qwen3-VL-4B-Instruct, and MiMo-VL-7B-SFT-2508 [coreteam2025mimovltechnicalreport].
+We additionally evaluate ADHint on Medical VQA to assess cross-domain generalization, using the Qwen2.5-VL-7B-Instruct backbone.
+For LLM math reasoning, we use Qwen2.5-Math-7B [yang2024qwen2], following the data and experimental settings of GHPO [liu2025ghpo].
+Training data.
+For MLLMs, we construct our training data from the ViRL39K dataset [wang2025vl], which spans multiple domains, including STEM, chart reasoning, document reasoning, and spatial reasoning.
+We use Qwen3-VL-30B-A3B-Thinking and Qwen3-VL-235B-A22B-Thinking to sample hints, and remove examples that are considered excessively simple for Qwen2.5-VL-7B-Instruct, resulting in a final set of 17K training queries, termed ViRL-Hint17k.
+For Medical VQA, we use a 15K subset of the PMC-VQA training set [zhang2023pmc].
+For LLMs, we use NuminaMath-S, a mixture of MATH [hendrycks2measuring] and NuminaMath-1.5 [hendrycksmath2021] curated by GHPO.
+More details are provided in Appendix D.1.
+Benchmarks.
+For MLLMs, we use benchmarks spanning diverse domains, including MathVista [lu2023mathvista], MathVerse [zhang2024mathverse], and We-Math [qiao2024we] for math reasoning; MMMU [yue2024mmmu] and MMMU-Pro [yue2024mmmupro] for multi-discipline reasoning; and LogicVista [xiao2024logicvista] for logical reasoning.
+For Medical VQA, we use the PMC-VQA test set (PMC-Test) [zhang2023pmc].
+For LLMs, we evaluate on math reasoning benchmarks, including AIME [patel2024aime], AMC [li2024numinamath], Minerva [lewkowycz2022solving], OlympiadBench [he2024olympiadbench], and MATH500 [hendrycks2measuring].
+More details are provided in Appendix D.2.
+Evaluation Metrics.
+For MLLMs, we report pass@1 and avg@8 for each benchmark [wang2025vl, wang2025perception].
+Pass@1 measures the accuracy of a single rollout per sample under greedy decoding, reflecting the model’s out-of-distribution reasoning generalization.
+Avg@8 computes the average accuracy over eight rollouts per sample at temperature 1.0, reflecting the model’s robustness and mastery of the underlying knowledge.
+For Medical VQA, we report pass@1 [zhu2025toward].
+For LLMs, we report either pass@1 or avg@32 depending on the benchmark to mitigate potential variance introduced by the small sizes of some datasets [liu2025ghpo].
+Implementation Details.
+For MLLMs, we train for 250 steps (about 2 epochs) with a learning rate of $1\times 10^{-6}$.
+We set the rollout batch size to 128 and the maximum generation length to 8K.
+For LLMs, we train for 5 epochs with an initial learning rate of $1\times 10^{-6}$, decayed to zero using a cosine schedule with a warm-up phase covering 10% of the total global steps.
+We set the maximum generation length to 2K.
+We set $w_{\text{max}}$ to 0.2 for MLLMs and 0.4 for LLMs, and use $\alpha=0.5$ as the CGM threshold for both.
+All training is performed on 4 nodes, each equipped with 8 H100 GPUs.
+Further details are provided in Appendix D.3.
+Baselines.
+We compare ADHint with several baselines: (1) GRPO; (2) SFT, trained for two epochs on the full training set; (3) SFT $\rightarrow$ RL, which applies two epochs of SFT followed by GRPO; (4) LUFFY [yan2025learning], which performs on-policy and off-policy updates jointly at each step; (5) StepHint [zhang2025stephint], which mixes hint-rollouts with multiple hint ratios, naive-rollouts, and ground-truth trajectories into a group for update; (6) HintGRPO [Huang_2025_ICCV], which progressively increases the hint ratio for the hardest samples to obtain positive trajectories; and (7) GHPO [liu2025ghpo], which treats hints as part of the query and uses multiple hint ratios for the hardest samples.
+We note that several baselines crash during training, and for these methods, we report results from the last checkpoint available before the crash.
+More detailed information about the baselines and their reimplementation is provided in Appendix D.4 and Appendix D.5, respectively.
+5.2 Main Results
+Table 1: Comparative performance on Qwen2.5-VL models.
+For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
+The upper block uses Qwen2.5-VL-3B as the base MLLM, while the lower block uses Qwen2.5-VL-7B.
+The best and runner-up results are highlighted in bold and underlined, respectively.
 Method Math Reasoning Multi-Discipline Logical Avg.
@@
 […]
-Table 1: Performance comparison of ADHint and baselines on Qwen2.5-VL across six benchmarks spanning three domains.
-For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
-The upper block uses Qwen2.5-VL-3B as the base MLLM, while the lower block uses Qwen2.5-VL-7B.
-The best and runner-up results are highlighted in bold and underlined, respectively.
-5.1 Settings
-Training data.
-We conduct extensive experiments across both MLLMs and LLMs to validate ADHint’s effectiveness.
-For MLLMs, we construct our training data from the ViRL39K dataset [wang2025vl], which spans multiple domains, including STEM, chart reasoning, document reasoning, and spatial reasoning.
-We use Qwen3-VL-30B-A3B-Thinking and Qwen3-VL-235B-A22B-Thinking [bai2025qwen3vltechnicalreport] to sample hints, and remove examples that are considered excessively simple for Qwen2.5-VL-7B-Instruct [bai2025qwen2], resulting in a final set of 17K training queries, termed ViRL-Hint17k.
-For LLMs, we use 21K examples sampled from public training data [yan2025learning], with hints generated by DeepSeek-R1 [guo2025deepseek].
-More details are provided in Appendix C.1.
-Benchmarks.
-For MLLM evaluation, we use six benchmarks spanning diverse domains, including MathVista [lu2023mathvista], MathVerse [zhang2024mathverse], and We-Math [qiao2024we] for math reasoning; MMMU [yue2024mmmu] and MMMU-Pro [yue2024mmmupro] for multi-discipline reasoning; and LogicVista [xiao2024logicvista] for logical reasoning.
-For LLM evaluation, we use AIME24, AIME25, AMC [li2024numinamath], Minerva [lewkowycz2022solving], OlympiadBench [he2024olympiadbench], and MATH500 [hendrycks2measuring].
-Further details are provided in Appendix C.2.
-Evaluation Metrics.
-For MLLM evaluation, we report pass@1 and avg@8 for each benchmark.
-Pass@1 measures the accuracy of a single rollout per sample under greedy decoding, reflecting the model’s out-of-distribution reasoning generalization.
-Avg@8 computes the average accuracy over eight rollouts per sample at temperature 1.0, reflecting the model’s robustness and mastery of the underlying knowledge.
-For LLM evaluation, we report either pass@1 or avg@32 depending on the benchmark to mitigate potential variance introduced by the small sizes of some datasets.
-Implementation Details.
-For MLLM training, we use Qwen2.5-VL-7B-Instruct, Qwen2.5-VL-3B-Instruct, and Qwen3-VL-8B-Instruct, and train them for 250 steps (about 2 epochs).
-We set the learning rate to $1\times 10^{-6}$, the rollout batch size to 128, and the maximum generation length to 8K.
-For LLM training, we use Qwen2.5-Math-7B [yang2024qwen2].
-Across all experiments, we set $w_{\text{max}}=0.2$ and $w_{\text{min}}=0.0$, the sampling radius of the noise term $\sigma$ to $0.01$, and the threshold in CGM to $\alpha=0.5$.
-We perform all training on 4 nodes, each equipped with 8 H100 GPUs.
-Further details are provided in Appendix C.3.
-Baselines.
-We compare ADHint with several baselines: (1) GRPO; (2) SFT, trained for two epochs on the full training set; (3) SFT $\rightarrow$ RL, which applies two epochs of SFT followed by GRPO; (4) LUFFY [yan2025learning], which performs on-policy and off-policy updates jointly at each step; (5) StepHint [zhang2025stephint], which mixes hint-rollouts with multiple hint ratios, naive-rollouts, and ground-truth trajectories into a group for update; (6) HintGRPO [Huang_2025_ICCV], which progressively increases the hint ratio for the hardest samples to obtain positive trajectories; and (7) GHPO [liu2025ghpo], which treats hints as part of the query and uses multiple hint ratios for the hardest samples.
-We note that several baselines crash during training, and for these methods, we report results from the last checkpoint available before the crash.
-More detailed information about the baselines and their reimplementation is provided in Appendix C.4 and Appendix C.5, respectively.
-Bench. Metric GRPO HintGRPO ADHint $\Delta$
-MathVista p@1 74.6 64.9 74.0 0.6
-a@8 78.8 64.4 79.6 0.8
-MathVerse p@1 63.1 64.2 75.2 12.1
-a@8 79.0 68.7 80.1 1.1
-We-Math p@1 73.1 68.6 76.4 3.3
-a@8 82.6 71.7 85.6 3.0
-MMMU p@1 61.0 47.2 64.4 3.4
-a@8 71.7 55.5 70.1 1.6
-MMMU-Pro p@1 46.5 38.4 51.8 5.3
-a@8 57.9 46.0 57.9 0.0
-LogicVista p@1 49.4 43.1 56.3 6.9
-a@8 63.0 51.9 64.9 1.9
-Avg. p@1 61.3 54.4 66.4 5.1
-a@8 72.2 59.7 73.0 0.8
-Table 2: Performance comparison of ADHint and baselines on Qwen3-VL-8B-Instruct.
-We report p@1 (pass@1) and a@8 (avg@8) for each method.
-The last column shows the absolute performance differences of ADHint relative to GRPO, where green indicates improvement and red indicates degradation.
-Figure 6: Performance comparison of ADHint and GRPO on Qwen2.5-Math-7B. For AIME and AMC, we report avg@32 due to their small sizes, and we report pass@1 for the remaining benchmarks.
+Table 2: Math reasoning performance on Qwen2.5-Math-7B. We report avg@32 on AIME and pass@1 on all other benchmarks. Results are taken from GHPO [liu2025ghpo].
+Method AIME AMC Minerva Math Olympiad Avg.
+GRPO [shao2024deepseekmath] 27.0 62.5 34.6 81.0 44.8 50.0
+LUFFY [yan2025learning] 31.1 67.5 37.9 81.8 44.0 52.5
+StepHint [zhang2025stephint] 29.6 61.4 36.4 81.0 40.4 49.8
+HintGRPO [Huang_2025_ICCV] 32.8 71.1 35.3 83.0 44.7 53.4
+GHPO [liu2025ghpo] 32.0 70.0 38.2 82.2 45.3 53.5
+[…]
+Table 3: Medical VQA performance on PMC-Test using Qwen2.5-VL-7B.
+Method PMC-Test [zhang2023pmc]
+GRPO [shao2024deepseekmath] 53.4
+LUFFY [yan2025learning] 49.4
+StepHint [zhang2025stephint] 48.3
+HintGRPO [Huang_2025_ICCV] 20.9
+GHPO [liu2025ghpo] 52.9
+[…]
+SOTA performance with MLLMs.
+As detailed in Tables 5.2, 3, and 5, ADHint achieves consistent gains across multiple backbones, validating its robust generalization regarding model capabilities, scales, and families.
+For Qwen2.5-VL models, ADHint consistently outperforms GRPO, yielding overall pass@1/avg@8 improvements of 2.1%/2.0% at the 7B scale and 2.3%/2.1% at 3B.
+Against recent hint-based methods, it delivers larger gains of 6.8%/11.1% (7B) and 9.7%/10.5% (3B).
+Furthermore, for Qwen3-VL models, ADHint maintains substantial effectiveness, boosting pass@1/avg@8 by 5.1%/0.8% at 8B and 4.1%/3.0% at 4B over the baselines.
+Beyond the Qwen series, ADHint also achieves a 2.0%/2.2% improvement over the baselines on MiMo-VL-7B-SFT-2508.
+We observe that several baselines perform worse than vanilla GRPO, which is consistent with our motivation and the training dynamics observed.
+Methods that use fixed, high hint ratios (e.g., GHPO, HintGRPO) can introduce excessive guidance and yield extremely high training rewards (often >0.9>0.9), but tend to overfit to the hint distribution, resulting in limited generalization and can even degrade reasoning ability without hints.
+Additionally, several methods (e.g., StepHint, LUFFY) ignore the imbalance in intra-group advantage estimation and the sharp gradients induced by hint tokens, leading to unstable update signals and training collapse.
+Moreover, our ViRL-Hint17k is more challenging, featuring longer reasoning chains (3100 tokens on average) than the relatively short baselines’ data (typically a few hundred tokens, e.g., $\sim$230 in GHPO), which further exposes these shortcomings.
+To address these issues, ADHint introduces novel mechanisms for hint-ratio schedule, relative-advantage estimation, and hint gradient modulation, providing a feasible solution for hint-based on-policy learning on challenging multimodal tasks.
+A more in-depth comparative analysis of ADHint and the baselines is provided in Appendix E.
+Table 4: Comparative performance on Qwen3-VL models. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
+Method MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista Avg.
+[…]
+GRPO [shao2024deepseekmath] 73.6/77.9 56.1/75.9 67.9/82.1 38.1/49.5 39.4/51.2 46.2/60.8 53.6/66.2
+LUFFY [yan2025learning] 64.7/72.3 56.3/70.9 56.6/74.5 29.9/46.6 32.0/49.7 38.2/57.1 46.3/61.9
+StepHint [zhang2025stephint] 66.2/71.2 53.7/70.2 58.2/74.0 34.7/52.5 32.8/49.4 38.0/57.0 47.3/62.4
+HintGRPO [Huang_2025_ICCV] 67.1/71.9 60.3/72.8 64.6/77.6 42.0/52.2 40.1/49.4 48.0/57.2 53.7/63.5
+GHPO [liu2025ghpo] 70.4/75.9 54.9/70.5 69.5/76.4 41.2/48.5 39.8/51.1 48.4/59.0 54.0/63.6
+[…]
+[…]
+GRPO [shao2024deepseekmath] 74.6/78.8 63.1/79.0 73.1/82.6 61.0/71.7 46.5/57.9 49.4/63.0 61.3/72.2
+LUFFY [yan2025learning] 54.4/53.7 28.0/18.6 26.0/18.1 21.0/26.6 21.2/27.2 14.5/21.8 27.5/27.7
+StepHint [zhang2025stephint] 55.0/54.8 28.4/20.5 25.8/17.5 22.5/27.7 22.3/27.2 20.5/22.5 29.1/28.4
+HintGRPO [Huang_2025_ICCV] 64.9/64.4 64.2/68.7 68.6/71.7 47.2/55.5 38.4/46.0 43.1/51.9 54.4/59.7
+GHPO [liu2025ghpo] 64.9/63.2 39.7/37.7 49.9/48.8 38.8/38.5 40.4/40.0 45.1/43.6 46.5/45.3
+[…]
+Table 5: Comparative performance on MiMo-VL-7B-SFT-2508.
+Method MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista Avg.
+GRPO [shao2024deepseekmath] 76.8/77.6 75.9/77.3 78.6/79.7 64.5/65.9 50.6/51.2 63.5/60.0 68.3/68.6
+LUFFY [yan2025learning] 70.2/68.1 64.5/68.7 60.7/67.5 48.2/61.4 35.2/47.4 41.7/53.7 53.4/61.1
+StepHint [zhang2025stephint] 72.3/67.9 64.5/67.8 63.2/66.6 36.2/43.3 37.0/47.2 43.3/54.6 52.8/57.9
+HintGRPO [Huang_2025_ICCV] 80.4/79.0 79.0/77.8 81.4/79.9 57.6/56.6 53.0/52.1 59.7/60.5 68.5/67.7
+GHPO [liu2025ghpo] 78.8/76.4 72.7/61.7 78.7/74.6 63.7/60.4 50.8/46.6 61.8/56.4 67.8/62.7
+[…]
+Performance with LLM.
+Table 3 evaluates ADHint on text-only reasoning.
+ADHint improves average accuracy by 2.4% over baselines across math reasoning benchmarks, indicating that it can broaden model capability boundaries across different modalities.
+Notably, the gains are consistent across benchmarks of varying difficulty, suggesting that ADHint benefits both routine and harder reasoning problems.
+Generalization to different domains.
+Table 3 reports results on the Medical VQA domain, which is knowledge-intensive, more OOD for the backbone models, and substantially different from standard math reasoning.
+We observe that the baselines do not generalize well, whereas ADHint improves accuracy by 1.7% over GRPO.
+These results suggest that ADHint can be readily transferred to more challenging real-world domains.
+5.3 Training Dynamics
+Figure 5 presents the training dynamics of several key metrics.
+Figure 4(c) shows that ADHint consistently maintains rewards within a moderate regime, yielding low-variance update signals.
+Figures 4(d) and 4(e) indicate that ADHint maintains exploration within a healthy entropy range with more stable gradient updates.
+Figures 4(f) and 4(g) show that ADHint steadily increases response length, while the proportion of overlength responses does not increase, which usually indicates repetition or garbled outputs.
+These trends suggest that the policy model continuously acquires effective long-term thinking capabilities.
+Taken together, these observations indicate that ADHint stably and adaptively learns reasoning ability from off-policy hints.
+(c) Training reward
+(d) Policy entropy
+(e) Gradient norm
+(f) Response length
+(g) Clipping ratio of responses
+(h) Format reward
+Figure 5: Training dynamics of key metrics for ADHint and GRPO.
+“ADHint: Hint/Naive” reports metrics computed on hint- or naive-rollouts only, while “ADHint: Overall” aggregates both rollout types.
+ADHint maintains stable learning and yields concurrent improvements in reasoning generalization and knowledge expansion.
+5.4 Ablation Study
+Table 6 reports the ablation results of four critical components in ADHint.
+Specifically, “w/o AH-SDP” replaces our adaptive hint-ratio schedule with an annealing hint-ratio strategy, while “w/o AE-RDP” pools hint-rollouts and naive-rollouts into a single group for advantage estimation.
+Removing any component results in performance degradation on both pass@1 and avg@8, underscoring the importance of our design.
+The difficulty-aware design in the hint-ratio schedule and relative-advantage estimation ensures efficient learning that balances exploration and imitation.
+The gradient modulation mechanisms prevent biased and destructive updates, ensuring healthy policy learning.
+Table 6: Ablation study of proposed ADHint. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
 Method MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista Avg.
@@
 w/o Selective Masking 74.7/73.7 58.5/61.0 67.4/67.6 55.8/55.9 41.6/40.9 47.3/46.3 57.6/57.6
-Table 3: Ablation study of proposed ADHint. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
+Table 7: Effect of the maximum hint ratio $w_{\text{max}}$.
 Setting MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista Avg.
@@
 $w_{\text{max}}=0.6$ 73.8/72.4 56.9/59.7 65.5/67.7 54.9/55.5 39.3/40.2 46.9/47.5 56.2/57.2
-Table 4: Effect of the maximum hint ratio $w_{\text{max}}$. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
+Table 8: Effect of the threshold $\alpha$ in CGM.
 Setting MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista Avg.
@@
 $\alpha=0.75$ 76.0/75.1 53.4/62.8 66.3/68.6 55.2/56.3 40.3/41.2 47.5/47.4 56.5/58.6
-Table 5: Effect of the threshold $\alpha$ in CGM. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
-5.2 Main Results
-SOTA performance with Qwen2.5-VL.
-In Table 5, we compare the proposed ADHint against several baselines on Qwen2.5-VL-7B and Qwen2.5-VL-3B.
-ADHint delivers consistent performance gains at both model scales.
-At the 7B scale, ADHint improves pass@1 by 2.1% and avg@8 by 2.0% over GRPO.
-Relative to recent methods, it delivers larger gains of 6.8% in pass@1 and 11.1% in avg@8.
-ADHint also performs well at the 3B scale, improving pass@1 by 2.3% and avg@8 by 2.1% over GRPO, and by 9.7% and 10.5% over recent methods, respectively.
-These results confirm ADHint’s advanced multimodal reasoning ability and suggest that integrating difficulty into both the hint-ratio schedule and relative-advantage estimation provides a feasible solution for hint-based on-policy learning on challenging multimodal tasks.
-In Appendix D, we provide a more in-depth comparative analysis of the performance differences between ADHint and the baselines.
-Performance with Stronger Backbone.
-Table 2 presents a comparison between ADHint and several baselines on Qwen3-VL-8B-Instruct, a recent strong backbone that achieves state-of-the-art performance among models of similar scale across multiple multimodal reasoning benchmarks.
-Compared with GRPO, ADHint achieves competitive performance on most metrics, with absolute gains of 5.1% in pass@1 and 0.8% in avg@8.
-Notably, ADHint achieves gains of over 5.0% in pass@1 on MathVerse, MMMU-Pro, and LogicVista, which cover challenging problems from three different domains.
-These results highlight the considerable promise of ADHint for improving general multimodal capabilities.
-As future work, we will collect more challenging datasets tailored to the Qwen3-VL series and leverage hints from stronger models to further unlock ADHint’s potential.
-Performance with LLM.
-In Figure 6, we further evaluate the effectiveness of ADHint on text-only reasoning.
-ADHint improves the average accuracy by 2.3% over GRPO across the six math reasoning benchmarks.
-On OlympiadBench, which covers Olympiad-level questions, and AIME25, which consists of challenging problems from the 2025 AIME, ADHint surpasses GRPO by 5.7% and 2.4%, respectively.
-These results demonstrate the strong generalization of ADHint, which broadens the capability boundaries of models across different modalities.
-5.3 Training Dynamics
-Figure 5 presents the training dynamics of several key metrics.
-Figure 5(a) shows that ADHint consistently maintains rewards within a moderate regime, yielding low-variance update signals.
-Figures 5(b) and 5(c) indicate that ADHint maintains exploration within a healthy entropy range with more stable gradient updates.
-Figures 5(d) and 5(e) show that ADHint steadily increases response length, while the proportion of overlength responses does not increase, which usually indicates repetition or garbled outputs.
-These trends suggest that the policy model continuously acquires effective long-term thinking capabilities.
-Taken together, these observations indicate that ADHint stably and adaptively learns reasoning ability from off-policy hints.
-5.4 Ablation Study
-Table 3 reports the ablation results of four critical components in ADHint.
-Specifically, “w/o AH-SDP” replaces our adaptive hint-ratio schedule with an annealing hint-ratio strategy, while “w/o AE-RDP” pools hint-rollouts and naive-rollouts into a single group for advantage estimation.
-Removing any component results in performance degradation on both pass@1 and avg@8, underscoring the importance of our design.
-The difficulty-aware design in hint-ratio schedule and relative-advantage estimation ensures efficient learning that balances exploration and imitation.
-The gradient modulation mechanisms prevent biased and destructive updates, ensuring healthy policy learning.
 5.5 Further Analysis
@@
 Both increasing and decreasing the hint ratio lead to performance degradation.
-As shown in Figure 5(a), a 20% hint ratio is sufficient to keep rewards within the range of 0.5–0.7, providing stable learning signals at a moderate difficulty level.
+As shown in Figure 4(c), a 20% hint ratio is sufficient to keep rewards within the range of 0.5–0.7, providing stable learning signals at a moderate difficulty level.
 Further increasing the hint ratio yields excessively high rewards for hint-rollouts, which introduces larger update bias.
-These results suggest that the prefix segment of reasoning trajectories is crucial for guiding the policy model toward effective reasoning.
+These results suggest that the initial prefix of the reasoning trajectory is critical for steering the policy model toward effective reasoning.
 Hint-token gradients should be restricted to a reasonable range.
@@
 6 Conclusion
-In this work, we introduce ADHint, which explicitly incorporates difficulty into hint-based on-policy learning.
-Enhanced by critical components that schedule the hint ratio with respect to the sample difficulty prior, estimate relative advantages with respect to the rollout difficulty posterior, and modulate token-level gradients based on distribution consistency, ADHint surpasses existing methods of comparable scale across various benchmarks, showcasing advanced reasoning abilities and out-of-distribution generalization.
+In this work, we propose ADHint, which explicitly incorporates difficulty into hint-based on-policy learning.
+Enhanced by critical components that schedule the hint ratio with respect to the sample difficulty prior, estimate relative advantages with respect to the rollout difficulty posterior, and modulate token-level gradients based on distribution consistency, ADHint surpasses existing methods across diverse settings, showcasing advanced reasoning abilities and OOD generalization.
 Despite its impressive performance, ADHint’s scaling behavior on larger models with 32B or more parameters, as well as its generalization to tasks without verifiable answers, merits further investigation.
-[…]
-Supplementary Material
-Our supplementary material is organized as follows:
-  • Section A presents the detailed workflow of ADHint.
-  • Section B provides a detailed discussion of related work.
-  • Section C provides additional details on training data construction, implementation details, and the reimplementation of baselines.
-  • Section D offers an in-depth comparison between ADHint and the baselines, analyzing the sources of their performance differences.
-  • Section E presents several case studies that provide a more intuitive view of ADHint’s behavior.
-A Training Algorithm of ADHint
+A Overview
+Our appendix is organized as follows:
+  • Section B presents the detailed workflow of ADHint.
+  • Section C provides an extended discussion of related work.
+  • Section D offers additional details on training data construction, experimental settings, and the reimplementation of baselines.
+  • Section E conducts an in-depth comparison between ADHint and the baselines, analyzing the sources of their performance differences.
+  • Section F provides several case studies to offer a more intuitive view of ADHint’s behavior.
+B Training Algorithm of ADHint
 In Algorithm 1, we provide the detailed workflow of ADHint as a complement to the main text.
@@
      Compute continuation entropy by Eq. 4
-     Compute token-wise factors $k_{i,t}$ by Eq. 5
+     Compute token-wise gradient modulation factors $k_{i,t}$ by Eq. 5
     end for
@@
   return $\pi_{\theta}^{*}$
-B Detailed Related Work
+C Detailed Related Work
 Multimodal Reinforcement Learning
-Group Relative Policy Optimization (GRPO) [guo2025deepseek, shao2024deepseekmath] removes the critic network from PPO [schulman2017proximal] and estimates advantages through intra-group sampling.
+GRPO [guo2025deepseek, shao2024deepseekmath] removes the critic network from PPO [schulman2017proximal] and estimates advantages through intra-group sampling.
 This approach has become a cornerstone for the development of recent reasoning models [yang2025qwen3, zeng2025glm, team2025kimi, team2025longcat, ye2025limo, wang2025reinforcement, tan2025towards].
@@
 There are also approaches that design more precise reward functions for tasks such as object detection and semantic segmentation [liu2025visionreasoner, liu2025visual], as well as methods that select or curate training data to improve efficiency [li2025truth, liang2025modomodomultidomaindatamixtures].
-However, these advances essentially remain optimizations of on-policy RL, primarily unlocking the latent potential of pretrained models, and they often fail to expand the models’ fundamental capability boundary.
-Reinforcement Learning with Off-Policy Data
+However, these advances primarily unlock the latent potential of pretrained models, often failing to expand their fundamental capability boundaries.
+Reinforcement Learning with Off-Policy Hints
 Recent work has investigated how to incorporate off-policy data into RL-based post-training.
@@
 UFT [liu2025uft] and Prefix-RFT [huang2025blending] employ the same temporally annealed hint ratios for all samples.
+SEELE [li2025staying] introduces learnable parameters to predict instance-specific hint ratios in a multi-round training scheme.
 Other methods reserve hints only for the hardest samples and determine the hint ratio from a fixed, predefined set via uniform sampling [chen2025data], binary search [zhang2025bread], or progressive search [Huang_2025_ICCV, liu2025ghpo].
 StepHint [zhang2025stephint] samples rollouts under multiple hint ratios and merges the resulting hint-rollouts, naive-rollouts, and ground truth off-policy data into a single group for advantage estimation.
-However, these methods overlook difficulty as a key factor, leading to insufficient and unstable training.
+However, these methods often overlook difficulty as a key factor, leading to insufficient and unstable training.
 In this work, we propose ADHint, which explicitly incorporates the sample difficulty prior into the hint-ratio schedule and the rollout difficulty posterior into relative-advantage estimation.
 This difficulty-aware design stabilizes training, enhances exploration, and expands the model’s knowledge boundary when learning from off-policy hints.
-(a) Avg@8 distribution of Qwen2.5-VL-7B (base model) on the 34K filtered samples.
-(b) Distribution of samples for which Qwen3-VL-30B-A3B-Thinking fails to produce any correct hint among eight candidates.
-(c) Avg@8 distribution of Qwen3-VL-235B-A22B-Thinking on the remaining hard samples.
-(d) Hint-length distribution and base model accuracy after hint sampling.
-Figure 7: Accuracy and length statistics during ViRL-Hint17k construction. At each sampling stage, we independently generate eight reasoning trajectories with temperature 1.0.
-C Detailed Experimental Settings
-C.1 Training data
+D Detailed Experimental Settings
+D.1 Training data
 For MLLM training, we construct an off-policy hint dataset based on ViRL39K [wang2025vl], which covers STEM, chart reasoning, document reasoning, and spatial reasoning.
@@
 Finally, we remove samples that are excessively simple for Qwen2.5-VL-7B and obtain ViRL-Hint17k, which we will open-source.
-For LLM training, we sample 21K examples with a maximum hint length of 4K tokens from public datasets [yan2025learning].
-C.2 Benchmarks
+For Medical VQA, we similarly employ Qwen3-VL-30B-A3B-Thinking to sample hints from the PMC-VQA training set [zhang2023pmc], ultimately constructing a filtered subset of 15K samples.
+For LLM training, we utilize NuminaMath-S curated by GHPO [liu2025ghpo], which comprises 18.3K samples combined from MATH [hendrycks2measuring] and NuminaMath-1.5 [hendrycksmath2021].
+(a) Avg@8 distribution of Qwen2.5-VL-7B (base model) on the 34K filtered samples.
+(b) Distribution of samples for which Qwen3-VL-30B-A3B-Thinking fails to produce any correct hint among eight candidates.
+(c) Avg@8 distribution of Qwen3-VL-235B-A22B-Thinking on the remaining hard samples.
+(d) Hint-length distribution and base model accuracy after hint sampling.
+Figure 6: Accuracy and length statistics during ViRL-Hint17k construction. At each sampling stage, we independently generate eight reasoning trajectories with temperature 1.0.
+D.2 Benchmarks
 For MLLM evaluation, we employ six benchmarks that cover mathematical reasoning, multidisciplinary multimodal reasoning, and logical reasoning:
@@
   • We-Math — A large-scale benchmark that investigates human-like mathematical reasoning in MLLMs, decomposing problems into sub-problems and introducing multi-dimensional metrics to analyze knowledge acquisition, generalization, and rote memorization behaviours.
-We use its testmini split.
+We use the testmini split.
   • MMMU — A massive multi-discipline multimodal understanding benchmark that spans college-level subjects such as science, engineering, humanities, and business, and demands deliberate integration of visual and textual information for complex problem solving.
@@
   • LogicVista — A benchmark that targets integrated logical reasoning in visual environments, focusing on compositional, step-by-step reasoning over diagrams and structured visual scenes.
-For LLM evaluation, we employ six mathematical reasoning benchmarks that align with our training domain:
+For Medical VQA, we evaluate on the PMC-VQA test set, a comprehensive benchmark derived from PubMed Central that assesses a model’s ability to comprehend medical images and generate clinically grounded answers.
+For LLM evaluation, we employ five mathematical reasoning benchmarks that align with our training domain:
   • AIME24 — A dataset that contains 30 competition-style problems from the 2024 American Invitational Mathematics Examination (AIME), a well-known high-school contest featuring challenging short-answer problems.
-  • AIME25 — The 2025 counterpart of AIME24, consisting of 30 similarly challenging problems from the 2025 AIME, which serves as an independent but distribution-matched test set.
   • AMC — A benchmark constructed from the American Mathematics Competitions, covering a broad range of high-school mathematics topics and providing multiple-choice questions that emphasize conceptual understanding and problem-solving skills.
@@
   • OlympiadBench — An Olympiad-level benchmark that comprises thousands of mathematically and scientifically challenging problems, offering a rigorous testbed for models’ higher-order reasoning and abstraction abilities.
-C.3 Implementation Details
-As a supplement to the main text, Tables 6 and 7 summarize the training configs of the GRPO and SFT baselines, respectively, while Table 8 reports the additional configs of ADHint.
+D.3 Implementation Details
+As a supplement to the main text, Tables 9 and 10 summarize the training configs of the GRPO and SFT baselines, respectively.
+Table 11 reports the default configs for ADHint, with the exception of MiMo-VL-7B-SFT-2508, Qwen2.5-Math-7B, and Medical VQA (based on Qwen2.5-VL-7B), where the maximum hint ratio $w_{\text{max}}$ is set to 0.4.
 Note that both GRPO and ADHint use 16 rollouts to ensure a fair comparison.
-C.4 Baselines
+D.4 Baselines
 We compare ADHint against a diverse set of baselines to validate its effectiveness.
@@
 We note that several baselines exhibit instability and crash during training.
-We provide a detailed analysis in Section D.
+We provide a detailed analysis in Section E.
+Table 9: Training configs for GRPO baselines.
 Model Config Value
-Qwen2.5-VL & Qwen3-VL max prompt length 4096
+Qwen2.5-VL & Qwen3-VL & MiMo-VL-SFT max prompt length 4096
 max response length 8192
@@
 learning rate 1e-6
-global batchsize 128
-micro batchsize 64
+rollout batchsize 128
+global batchsize 64
 num rollouts 16
@@
 max pixels 128028281280*28*28
-Qwen2.5-Math max prompt length 1024
-max response length 4096
+Qwen2.5-Math max prompt length 2048
+max response length 2048
 temperature 1.0
 learning rate 1e-6
-global batchsize 128
-micro batchsize 64
-num rollouts 16
-epoch 1
-Table 6: Training configs for GRPO baselines.
+lr schedule cosine
+warmup ratio 0.1
+rollout batchsize 128
+global batchsize 32
+epoch 5
+Table 10: Training configs for SFT baselines.
 Method Config Value
@@
 epoch 2
-min pixels 2562828256*28*28
-max pixels 128028281280*28*28
-Table 7: Training configs for SFT baseline.
+min pixels $256\times 28\times 28$
+max pixels $1280\times 28\times 28$
+Table 11: Training configs for ADHint.
 Method Config Value
@@
 num of hint-rollouts $m$ 8
-minimum hint ratio $w_{\text{min}}$ 0.0
 maximum hint ratio $w_{\text{max}}$ 0.2
@@
 threshold of CGM $\alpha$ 0.5
-Table 8: Additional training configs for ADHint.
-C.5 Reimplementation Details of Baselines
+D.5 Reimplementation Details of Baselines
 For all baselines, we strictly follow the official implementations and the settings reported in the original papers.
@@
 Following the original paper, we omit the KL constraint and apply a 5-step GRPO warmup.
-D Comparative Analysis of ADHint and Baselines
-As shown in Table 5 and Figure 2, ADHint consistently outperforms all baselines.
+E Comparative Analysis of ADHint and Baselines
+As shown in Figure 2, ADHint consistently outperforms all baselines.
 Here, we provide a detailed comparative analysis of ADHint and these baselines to better understand the sources of their performance differences.
 SFT learns, RL generalizes.
-Compared with pure GRPO, the $\mathrm{SFT}\rightarrow\mathrm{RL}$ pipeline improves avg@8 by 1.3% but reduces pass@1 by 12.2% at the 7B scale.
+As shown in Table 5.2, compared with pure GRPO, the $\mathrm{SFT}\rightarrow\mathrm{RL}$ pipeline improves avg@8 by 1.3% but reduces pass@1 by 12.2% at the 7B scale.
 This pattern indicates that SFT injects new knowledge into the model and raises the upper bound of its performance when multiple samples are allowed, whereas the subsequent RL phase, starting from an SFT-initialized policy, tends to lose part of the exploration capability that is crucial for strong pass@1 performance compared with directly training with on-policy RL.
@@
 In contrast, HintGRPO trains more stably and maintains a high average reward during training.
-However, as shown in Table 5, HintGRPO performs poorly on the benchmarks and shows limited generalization.
+However, as shown in Table 5.2, HintGRPO performs poorly on the benchmarks and shows limited generalization.
 The progressive hinting strategy for the hardest samples helps obtain positive rollouts, but it ignores sample difficulty and can provide excessive guidance, causing SFT-like overfitting to the hint distribution and weakening exploration.
@@
 Imbalanced intra-group advantage estimation leads to abnormal updates.
-StepHint [zhang2025stephint] mixes hint-rollouts with multiple fixed hint ratios, naive-rollouts, and ground-truth trajectories into a single group for advantage estimation without distinguishing among them.
+Methods like StepHint [zhang2025stephint] mix hint-rollouts with multiple fixed hint ratios, naive-rollouts, and ground-truth trajectories into a single group for advantage estimation without distinguishing among them.
 As shown in Figure 1(b) of the main text, this imbalanced intra-group estimation allows longer hint-rollouts with more positive trajectories to receive disproportionately large advantages, which biases policy updates toward imitating hints and leads to abnormal update behavior.
@@
 CGM and Selective Masking further modulate token-level gradients within hints, which promotes more stable and reliable policy updates.
-E Case Study
+F Case Study
 Figure 7 shows a representative training example.
-This case includes the extra prompt we add beyond the user query, which instructs the model to think before producing its response.
+This case includes the extra prompt we added beyond the user query, which instructs the model to think before producing its response.
 For clarity, we omit this additional prompt in the subsequent cases.
@@
 However, when we give the policy a hint taken from the complete off-policy reasoning trajectory, it continues the reasoning process and computes the final answer $z=2\sqrt{15}$.
-In this way, the model not only learns the essential knowledge required to solve the problem from the hint, but also improves its reasoning ability.
+In this way, the model not only learns the essential knowledge required to solve the problem from the hint but also improves its reasoning ability.
 Figure 8 shows two representative benchmark examples.